### v1

Abstract
AI methods, such as generative models and reinforcement learning, have recently been applied to combinatorial optimization (CO) problems, especially NP-hard ones. This paper compares such GPU-based methods with classical CPU-based methods on Maximum Independent Set (MIS). Experiments on standard graph families show that AI-based algorithms fail to outperform and, in many cases, to match the solution quality of the state-of-art classical solver KaMIS running on a single CPU. Some GPU-based methods even perform similarly to the simplest heuristic, degree-based greedy. Even with post-processing techniques like local search, AI-based methods still perform worse than CPU-based solvers.
We develop a new mode of analysis to reveal that non-backtracking AI methods, e.g. LTFT (which is based on GFlowNets), end up reasoning similarly to the simplest degree-based greedy approach, and thus worse than KaMIS. We also find that CPU-based algorithms, notably KaMIS, have strong performance on sparse random graphs, which appears to refute a well-known conjectured upper bound for efficient algorithms from Coja-Oghlan & Efthymiou (2015).
1 Introduction
Combinatorial optimization (CO) lies at the core of numerous scientific and engineering studies, encompassing applications in network design, resource allocation, healthcare, and supply chain (Du & Pardalos, 1998; Hoffman, 2000; Zhong & Tang, 2021). Combinatorial optimization usually involves selecting an optimal solution from a discrete but often exponentially large set of candidates. Many are NP-hard (meaning that if $P\neq NP$ then there is no polynomial-time algorithm for solving them in the general cases (Papadimitriou & Steiglitz, 1998)). This makes it challenging to design algorithms with provable guarantees, but in practice solvers can find reasonable quality solutions (e.g., Gurobi (Gurobi Optimization, LLC, 2024)).
Recent advances in artificial intelligence (AI) and GPU computing have motivated use of AI-inspired approaches, e.g., Graph Neural Networs (GNNs) and reinforcement learning, to learn problem-specific strategies for NP-hard optimization problems such as MIS (Li et al., 2018; Ahn et al., 2020) and TSP (Kool et al., 2018; Zhang et al., 2021). AI models can also be trained to predict search directions or refine heuristic rules (Li et al., 2018; d O Costa et al., 2020). These algorithms utilize advanced GPUs and often require shorter inference compared to CPU-based algorithms. Additionally, AI-inspired methods avoid the need to design heuristics for specific problems, allowing generalization to new instances and problems (Bengio et al., 2021b; Cappart et al., 2023).
Despite the claimed benefits of AI-inspired methods, a few years ago Angelini & Ricci-Tersenghi (2023) showed that a Graph Neural Network based MIS algorithm (Schuetz et al., 2022) failed to surpass greedy algorithms. Böther et al. (2022) showed that AI-inspired approaches fail to provide superior search directions compared to traditional heuristics in tree search algorithms for MIS. Gamarnik (2023) suggests that GNN has theoretical limits which may become obstacles for GNN-based MIS algorithms.
However, in recent years many new AI-inspired CO algorithms, utilizing a variety of techniques, such as diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024), GPU-accelerated sampling (Sun et al., 2023), and direct optimization (Alkhouri et al., 2024) have been developed and claimed to significant improve over the previous ones. Furthermore, GFlowNets Bengio et al. (2021a), which have been proposed as general-purpose tools for tasks like scientific discovery and reinforcement learning Bengio et al. (2023), has also been used to solve CO problems (Zhang et al., 2023). Since combinatorial optimization is an arena where humans have hand-designed algorithms for many decades, the following question is of great scientific interest:
Do AI-inspired algorithms perform better than classical heuristics for
combinatorial optimization?
1.1 Our contributions
We explore the question in the context of Maximum Independent Set (MIS) problem: given a graph, aiming to find the largest subset of nodes with no edges present between any node pair. The simplicity of the problem attracted design of many heuristics to tackle the problem (Andrade et al., 2012; Lamm et al., 2015). In recent works, MIS is also a main target in efforts that design AI-inspired approaches such as non-convex optimization (Schuetz et al., 2022; Alkhouri et al., 2024), reinforcement learning (Ahn et al., 2020; Zhang et al., 2023), and diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024).
For classical heuristics, we test degree-based greedy (Deg-Greedy, pick a node with smallest degree at each step), and the state-of-the-art MIS solver KaMIS (Lamm et al., 2017; Dahlum et al., 2016; Hespe et al., 2019). For AI-inspired algorithms, we test the newest algorithms from each “category” according to the techniques they use, including sampling algorithm iSCO (Sun et al., 2023), non-convex optimization algorithm PCQO (Alkhouri et al., 2024), reinforcement-learning related algorithms LwD (Ahn et al., 2020) and LTFT (Zhang et al., 2023) (based on GFlowNets (Bengio et al., 2021a)) and diffusion models DIFUSCO (Sun & Yang, 2023) and DiffUCO (Sanokowski et al., 2024).
Testing on different graph types with different sizes and densities leads to the following empirical finding (Section 3).
Finding 1: Current AI-inspired algorithms for MIS still don’t outperform the best heuristic KaMIS, which runs on a single thread in a CPU, while AI-inspired methods often require significant computational resources.
Finding 2: As the graph becomes larger or denser, KaMIS exhibits a notable superiority to AI-inspired algorithms.
Finding 3: The simplest degree-based greedy algorithm (Deg-Greedy) serves as a very strong baseline. Some AI-inspired algorithms perform similarly to or worse than Deg-Greedy, especially for larger and denser graphs.
Section 4 presents ablation studies to understand why some AI-inspired methods fail to improve over the simplest Deg-Greedy method. We introduce a new mode of analysis, serialization, that transforms the solution of any algorithm into a sequential list of choices leading to the final independent set. We compare sequential order with the one produced by Deg-Greedy (Sections 4.1 and 4.2). We find that the reinforcement learning related algorithm LTFT, based on GFlowNets, behaves similarly to Deg-Greedy. We also found several qualitative characteristics that appear to distinguish algorithms that perform significantly better than Deg-Greedy from those that perform similarly or even worse than Deg-Greedy. We also explore whether AI-inspired method can improve their solution quality via a post-processing step using local search (Section 4.3), but find that they still fail to outperform KaMIS after some improvements. In Section 4.4, we discuss an additional result that may be of interest for MIS experts: on random graphs, KaMIS has a level of performance on MIS that appears to contradict a well-known conjecture about polynomial-time algorithms Coja-Oghlan & Efthymiou (2015).
2 Benchmarking MIS Algorithms
We focus the experiment setup for benchmarking different algorithms for Maximum Independent Set problems (MIS).
2.1 Maximum Independent Set (MIS) problem
Given an undirected graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ where ${\mathcal{V}}$ is the set of nodes and ${\mathcal{E}}$ is the set of edges, an independent set is a subset of vertices ${\mathcal{I}}\subseteq{\mathcal{V}}$ such that no two nodes in ${\mathcal{I}}$ are adjacent, i.e., $(u,v)\notin{\mathcal{E}}$ for all $u,v\in{\mathcal{I}}$. The goal in MIS is to find the largest possible independent set, ${\mathcal{I}}^{*}$.
2.2 MIS algorithms
We classify the algorithms we test as: (1) classical heuristics, which includes Deg-Greedy and KaMIS (OnlineMIS and ReduMIS); (2) GPU-accelerated non-learning algorithms, which includes iSCO and PCQO; and (3) learning-based algorithms, which includes LwD, LTFT, DIFUSCO, and DiffUCO.
Deg-Greedy
(Degree-based greedy) Simplest heuristic: always picks a node with the smallest degree in the current graph, add to the current independent set, and delete that node and all of its neighbors from the graph. Most papers on AI-inspired methods do not compare with this baseline.
OnlineMIS and ReduMIS
are two variants of the MIS solver KaMIS, mainly consists of three alternating steps: greedy, local search, and graph reductions. OnlineMIS (Dahlum et al., 2016) only applies a simple reduction after local search, while ReduMIS Lamm et al. (2017) applies many graph reduction techniques.
iSCO
(Sun et al., 2023) is a GPU-accelerated sampling-based method, incorporating gradient-based discrete MCMC and simulated annealing. The MCMC is designed based on the Metropolis-Hasting algorithm, which if given enough time (exponential), can get the optimal solution.
PCQO
(Alkhouri et al., 2024) directly optimizes the quadratic loss function of the MIS using gradient descent. It is sensitive to optimization hyperparameters, so hyperparameter search is required for achieving good results.
LwD
(Ahn et al., 2020) is a reinforcement learning based algorithm which models the problem as a Markov Decision Process (MDP) and requires a dataset (without solutions) to train the policy. In each step, several nodes are added to the independent set and are never deleted. We call it a non-backtracking algorithm. Ahn et al. (2020) reported that it outperforms KaMIS on very large but very sparse random graphs, which we do not include in our benchmark.
LTFT
(Zhang et al., 2023) is also a non-backtracking MDP-based algorithm similar to LwD, but it only selects one node at a time, which is decided by GFlowNets (Bengio et al., 2021a). Thus, it has a very similar procedure to Deg-Greedy. The algorithm requires a dataset (without solutions) to train the neural network.
DIFUSCO and DiffUCO
DIFUSCO (Sun & Yang, 2023) is an end-to-end one-shot MIS solver using diffusion models and requires a supervised training dataset with solutions. DiffUCO (Sanokowski et al., 2024) also uses diffusion model but with unsupervised learning and annealing techniques. It requires a training dataset without solutions.
Non-backtracking vs one-shot Among algorithms that build up the set step by step, Deg-Greedy, LTFT and LwD are non-backtracking, meaning once a node is added to the set it is never dropped from it. OnlineMIS and ReduMIS are backtracking algorithms, since as part of local search they might decide that a previously picked should be dropped from the set to allow further additions. AI-inspired methos PCQO, DIFUSCO, and DiffUCO are one-shot algorithms, since they work like end-to-end MIS solvers and directly return the full set.
[…]
CPU-based GPU-acc Learning-based $n$ $d$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 10 29.25 30.50 30.50 30.62 25.00 30.38 28.62 30.25 30.02 30 13.63 14.00 14.75 14.50 11.50 14.38 13.12 13.88 13.92 300 10 77.50 93.88 94.38 94.75 90.62 94.25 88.62 93.50 91.84 30 44.50 47.88 47.88 47.62 41.75 46.88 43.25 43.88 45.04 100 16.13 18.00 18.38 18.00 13.38 17.00 16.25 16.62 16.93 1000 10 303.25 314.75 316.13 315.62 305.12 311.25 297.00 303.88 311.67 30 151.00 158.88 163.75 163.50 143.00 158.38 150.00 143.75 152.55† 100 60.63 64.75 66.63 66.50 55.62 63.88 60.88 55.38 63.55† 300 22.25 25.00 25.75 24.62 20.00 19.12* 22.62 20.88 18.36† 3000 10 907.13 947.25 954.25 950.88 919.00 934.12 888.25 902.00 935.28† 30 451.88 480.88 493.13 491.62 442.38 473.25 449.00 413.38 459.57† 100 183.63 194.38 201.50 200.38 170.38 190.75* 184.00 171.38 194.07† 300 73.50 77.63 80.75 78.88 63.50 73.88 1000 23.38 26.00 26.25 19.88 23.62 10000 10 2999.88 3161.88 3173.62 3149.92 2685.62 30 1498.00 1607.50 1639.88 1625.47 1334.75 100 613.75 650.00 670.88 573.63 300 249.00 258.38 272.25 239.13 1000 87.13 91.50 94.25 80.88 3000 29.63 33.00 33.25 19.00
2.3 Graph types
Erdős-Renyí (ER) graph
(Erdös & Rényi, 1959) are random graphs where edges are connected uniformly at random (with a given probability or a fixed number of edges). We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing the number of edges at $\frac{nd}{2}$. Previous benchmark (Böther et al., 2022) and algorithms (Ahn et al., 2020; Sun & Yang, 2023; Zhang et al., 2023; Alkhouri et al., 2024) used it as test graphs for MIS, though without varying parameters as we did.
Barabási–Albert (BA) graph
(Albert & Barabási, 2002) are random graphs generated by a probabilistic growth process, mimicking real-world networks such as Internet, citation networks, and social networks (Albert & Barabási, 2002; Radicchi et al., 2011). For BA graphs, we vary $2$ parameters: number of nodes $n$ and parameter $m$ (not number of edges). The average degree of BA graphs can be approximated as $2m$.
Real-world graphs
We pick REDDIT-MULTI-5K and COLLAB (Yanardag & Vishwanathan, 2015) from TUDataset website (Morris et al., 2020), since they have enough graphs for training and graph sizes not too small. REDDIT-MULTI-5K has $508.52$ average nodes and $594.87$ average edges. They are mostly very sparse graphs. COLLAB has $74.49$ average nodes and 2457.78 average edges. They are mostly small but dense graphs.
2.4 More experiment details
For synthetic graphs, we test $8$ graphs for each parameter $(n,d)$ or $(n,m)$. We test on $100$ graphs for real-world datasets. For learning-based algorithms, we use $4000$ training graphs generated using the same parameter (in case of random graphs) or drawn from the same real-world dataset. For algorithms requiring hyperparameters, we use default hyperparameters in most cases (Details in Appendix B).
We set $24$-hr time limit for KaMIS (OnlineMIS and ReduMIS) since it runs on a single CPU thread with 32GB memory, and our benchmark focuses on performance on solution quality. Note that AI-based methods run well only on relatively small graphs, and ReduMIS runs in less than one hour on small graphs $(n\leq 3000)$. Given 24 hours ReduMIS can handle much larger graphs (up to $n\approx 1e6$).
For iSCO and learning-based algorithms, we report results within our computational limit (generally a single 80GB A100 for $96$hrs, details in Appendix B.) We test PCQO for $n\leq 10000$ because the performance degrades quickly for large graphs using default hyperparameter search domain.
3 Experiment Results and Main Findings
In this section, we present our main experiment results. The performance of different algorithms on Erdős-Renyí (ER) graphs, Barabási–Albert (BA) graphs, and real-world graphs are shown in Tables 1, 2 and 3 respectively.
AI-inspired algorithms don’t outperform ReduMIS.
Our first main finding is that, current AI-inspired algorithms do not outperform the best classical heuristics ReduMIS in terms of performance. As shown in Tables 1, 2 and 3, ReduMIS consistently achieves superior results compared to all other methods, with the exception of iSCO sometimes perform similarly, on ER, BA, and real-world graphs.
Additionally, although learning-based algorithms are claimed to be more efficient, they require significant training time and GPU memory, which is over our resource constraint for graphs with more than 3000 nodes, while ReduMIS can handle graphs with up to $1\times 10^{6}$ nodes (See Table 7). Although iSCO performs close to ReduMIS, it requires significant GPU memory and we are unable to get results for dense graphs with 10000 nodes. Its performance also become worse than ReduMIS for larger graphs.
We also note that LwD performs the best among learning-based algorithms, despite it being the oldest learning-based algorithm we tested. In summary, our experiment results show that current AI-inspired algorithms still don’t outperform the best classical heuristics for the MIS problem.
[…]
Heuristics GPU-acc Learning-based $n$ $m$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 5 39.50 39.50 39.50 39.50 33.88 39.50 38.12 39.38 39.25 15 21.00 21.63 21.63 21.62 18.38 21.62 20.62 21.25 21.25 300 5 121.88 123.13 123.13 123.12 117.62 123.12 115.62 122.88 123.00 15 69.38 71.38 71.38 71.38 65.62 70.75 64.75 69.50 69.25 50 38.00 49.88 50.00 50.00 27.25 50.00 41.75 50.00 50.00 1000 5 412.63 417.13 417.13 417.12 410.88 416.00 385.75 417.12 416.00 15 232.75 245.00 246.38 246.25 231.00 243.12 224.25 236.38 240.75 50 107.50 115.75 116.88 116.75 107.88 113.00* 106.38 105.25 56.38† 150 81.50 150.00 150.00 150.00 109.62 150.00* 82.25 135.00† 3000 5 1236.63 1257.00 1257.13 1255.62 1238.75 1248.12 1177.25 1254.38 1252.63† 15 709.50 749.63 754.50 752.00 710.62 727.75* 661.88 726.38 737.38† 50 335.38 362.63 369.75 368.25 338.88 336.88* 323.88 142.25† 150 147.25 160.25 165.75 164.00 144.00 144.75 500 141.00 500.00 500.00 500.00 223.88 10000 5 4128.75 4206.00 4205.38 4190.47 4161.88 15 2377.88 2525.00 2534.00 2517.07 2428.13 50 1122.13 1228.88 1251.13 1241.98 1184.63 150 511.25 555.75 580.13 573.37 537.63 500 192.50 209.13 217.50 198.75 1500 67.25 1500.00 1500.00 1500.00
Table 3: (Comparison of the performance of different algorithms on real-world graphs) Table 1 on real-world datasets. In general, graphs in REDDIT-MULTI-5K are very sparse, while COLLAB are dense but small. The numbers within ±$\pm 1\%$ of the best are highlighted.
Dataset Heuristics GPU-acc Learning-based Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO REDDIT-MULTI-5K 350.73 350.73 350.66 350.73 344.47 350.73 343.35 350.72 350.69 COLLAB 8.68 8.70 8.70 8.70 8.57 8.69 8.70 8.75 8.70
The performance gap between ReduMIS and AI-based methods widens with larger and denser graphs.
While ReduMIS consistently outperforms AI-based methods in most cases, the performance gap is small on small or sparse graphs. On ER graphs when $n=100,300$ and average degree $d=10$, LwD has results within $1\%$ gap from ReduMIS. On BA graphs for $n=300,1000$ and $m=5$, LwD, DIFUSCO, and DiffUCO has less than $0.5\%$ gap from ReduMIS. Across real-world datasets all tested algorithms produce independent sets of similar sizes on average. However, as the graph becomes larger or denser, the performance gap between ReduMIS and AI-based algorithms enlarges. On ER graphs, when $n=1000$, there is a clear performance gap between AI-based algorithms and ReduMIS, with the only exception of the sampling based iSCO. When $d=10$, DiffUCO and LwD still reaches 98% of ReduMIS’s performance. For denser graphs($d=100$), DiffUCO and LwD only reaches 96% of ReduMIS’s performance. This gap widens further for $n=3000$, where AI-based algorithms perform significantly worse than ReduMIS and sometimes fail to outperform simple heuristic Deg-Greedy. A similar trend has also been observed on BA graphs. Classical solvers have no difficulty handling graphs with a million edges, but learning-based implementations struggle to scale up to that size.
Deg-Greedy serves as a strong baseline.
Another key finding is that the simplest degree-based greedy (Deg-Greedy) serves as a remarkably strong baseline. As shown in Tables 1, 2 and 3, leveraging neural networks for node selection, LTFT often perform comparably to Deg-Greedy, particularly on larger or denser graphs. For example on ER graphs when $n=1000$ or $3000$, LTFT gives performance within 2% of Deg-Greedy (Table 1). An exception is for BA graphs where the graph is very dense ($n=3000,m=500$) and clear patterns emerge in the graph, LTFT surpasses Deg-Greedy. Additionally, DIFUSCO and PCQO fail to outperform Deg-Greedy on larger or less sparse graphs, such as $n=3000$, $d>30$.
4 AI-Inspired vs Deg-Greedy
Our results show that the state-of-the-art AI-inspired algorithms for MIS still do not outperform the best heuristic ReduMIS. The surprising finding was that they also often do not outperform the simplest classical heuristic, Deg-Greedy, especially on large and dense graphs. In this section, we delve deeper into this comparison (Sections 4.1 and 4.2). Furthermore, we explore the impact of augmenting various algorithms with a local search as a post hoc step to enhance solution quality (Section 4.3).
4.1 Comparison between Deg-Greedy and LTFT
Figure 1: Percentage of rounds when LTFT selects the node with smallest possible degree, i.e., behaves similarly to degree-based greedy. On ER graphs when the graph is dense (closer to the bottom left corner), the percentage of rounds LTFT selects the node with smallest possible degree is higher, i.e., behaves more similarly to degree-based greedy. On BA graphs, the percentage to choose the smallest possible node is generally lower, but on denser BA graphs LTFT also behaves more like degree-based greedy.
Deg-Greedy sequentially picks nodes for the independent set. At each step, it picks the node with the smallest degree in the residual graph (where the nodes in the independent set and their neighbors are removed). It does not reverse any decisions (ie once picked, the node stays in the independent set). As in Section 2.2 we call it a non-backtracking algorithm. LTFT is also a non-backtracking algorithm and it often perform similarly to Deg-Greedy in Tables 1 and 2. It uses a trained policy network GFlowNets (Bengio et al., 2021a) to pick a node for the independent set at each step. Thus, we can naturally compare LTFT with Deg-Greedy by investigating how close this trained policy compares to the naive policy in Deg-Greedy.
The results are shown in Figure 1. Overall, we observe that LTFT frequently selects nodes with very small degrees. On ER graphs where the average degree is at least 30, LTFT picks the node with the smallest degree over 85% of rounds except for sparse graphs ($d=10$ and $(n,d)=(3000,30)$). On BA graphs, while the percentage is lower, it still exceeds 75% on large and sufficiently dense graphs. Moreover, in cases where LTFT selects the smallest degree nodes less frequently, its performance is worse than Deg-Greedy.
To conclude, our results suggest that despite using neural net to learn the policy, LTFT is closely aligned with Deg-Greedy: prioritizing nodes with small degrees at each step. The high consistency between the node selection strategies of LTFT and Deg-Greedy can explain their similar performance.
4.2 Serialization: allows comparing to Deg-Greedy
Figure 2: The percentage to choose the smallest possible degree node on different part of the (degree-based) serialization. We find that the best algorithms (OnlineMIS, ReduMIS, iSCO) and the best performing learning-based algorithm LwD share a similar characteristic pattern, that they have high consistency with degree-based greedy on second and third part of the serialization, while on the first part there is a low consistency. On the other hand, PCQO and DIFUSCO has low consistency with degree-based greedy in general.
In the previous part, we demonstrate that LTFT employs a heuristic similar to Deg-Greedy, selecting the node with the smallest degree in the remaining graph in each round. However, other algorithms, such as ReduMIS, PCQO, and DIFUSCO, which does not select one node at a step without backtracking, cannot be analyzed based on the sequence of nodes picked. Thus, we introduce a method called degree-based solution serialization to analyze their behavior and compare them with Deg-Greedy.
Given a graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ and an independent set solution ${\mathcal{I}}$ (which is an independent set), process proceeds as follows: (1) Repeatedly remove the node in ${\mathcal{I}}$ with the smallest degree. (2) After removing a node from ${\mathcal{I}}$, also remove it and its neighbors in ${\mathcal{G}}$. (3) Continue this process until all nodes in ${\mathcal{I}}$ are removed. The order in which nodes are removed forms the serialization of the solution. This procedure is detailed in Algorithm 1.
To evaluate the algorithms, we compute the percentage of rounds in which the smallest degree node is selected during serialization, similar to our comparison between LTFT and Deg-Greedy. Due to random tie-breaking in Algorithm 1, we repeat the process 100 times and select the serialization with the highest number of rounds selecting the smallest degree node. Although Deg-Greedy theoretically achieves 100% smallest-degree selections in its best serialization, random tie-breaking prevents us from perfectly recovering this with 100 repetitions. Instead of reporting the overall percentage, we divide the serialization into three equal parts and report the percentages for each part.
The results, shown in Figure 2, compare the percentage of smallest-degree node selections across different algorithms. Due to space constraints, we present results only for ER and BA graphs under selected parameters; additional results are available in Section C.2.
Algorithms that often perform the best, namely OnlineMIS, ReduMIS, and iSCO, exhibit a consistent pattern after serialization. In the first one-third of the serialization, these algorithms deviate significantly from selecting the smallest-degree nodes. However, in the middle and final thirds, the percentage of smallest-degree node selections increases substantially. This suggests that while degree-based greedy heuristics may appear shortsighted initially, they are highly effective in the later stages of solution construction. Interestingly, LwD, which performs the best among learning-based methods tested in our setting, also shares this pattern.
As for PCQO and DIFUSCO, they show consistently low percentages of smallest-degree node selections throughout the serialization, particularly on ER graphs.
Through serialization, we observe distinct differences in node selection patterns among algorithms. Our findings suggest that AI-based methods might fail to utilize (PCQO and DIFUSCO) or emphasize too much (LTFT) on simple yet highly effective heuristics, such as greedily selecting the smallest-degree node, which may partly explain their performance limitations.
In Section C.3, we perform a pseudo-natural serialization for LwD. LwD is also a non-backtracking algorithm like LTFT, but it does not have a “natural serialization” like LTFT (Section 4.1) because it chooses several nodes in a step. The pseudo-natural serialization performs serialization in each step of LwD. The results in Figure 6 align with our “counterfactual” serialization results here.
4.3 Incorporating local search to improve solution
In the previous sections, we show that solutions generated by AI-based algorithms generally differ from those produced by degree-based greedy methods, which may explain their inferior performance on MIS problems. A natural idea is to enhance these solutions with simple heuristics, such as local search. Local search post-processing has also been used for AI-algorithms in previous works (Ahn et al., 2020; Böther et al., 2022).
We applied the $2$-improvement local search (Andrade et al., 2012) (details in B.7), which is used in KaMIS, as a post-processing step to all algorithms (except OnlineMIS and ReduMIS since they already has local search), and the resulting performance improvements are presented in Table 4.
Table 4: Incoperating local search as a post-processing procedure We add $2$-improvement local search for the algorithm tested. We report the performance after the local search and the improvement from local search in (). AI-inspired methods still don’t outperform ReduMIS after local search. We also observe that if the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search. Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search.
Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 148.50 (5.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 59.00 (3.38) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 453.88 (11.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) BA 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 235.50 (4.50) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 110.25 (2.37) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 723.50 (12.88) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37)
As shown in Table 4, algorithms like PCQO and DIFUSCO benefit significantly more from the local search post-processing compared to others, such as Deg-Greedy and LwD. This observation aligns with our earlier findings: If the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search. Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search.
In addition, although all algorithms except iSCO have improvements after local search, they still perform worse than ReduMIS in most cases.
In summary, our analysis highlights a promising direction for designing machine learning-based combinatorial optimization algorithms. Rather than relying solely on end-to-end methods like PCQO or DIFUSCO, incorporating classical heuristics, such as greedily selecting the smallest-degree node, into the overall algorithm may yield better results. One potential approach could involve using machine learning algorithms to identify a small subset of nodes, followed by a degree-based greedy method to complete the solution.
4.4 KaMIS refutes famous conjecture?
For theoretical analysis for MIS on ER graphs and regular graphs, see Coja-Oghlan & Efthymiou (2015); Wormald et al. (1999); Barbier et al. (2013); Gamarnik & Sudan (2014). In ER graphs with $n$ nodes and average degree $d$, the MIS has size $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$, and simplest random greedy achieves half-optimal at $\frac{n\ln d}{d}$ (Grimmett & McDiarmid, 1975). Yet, there is no known polynomial-time algorithm which can achieve MIS size $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ and it is conjectured that polynomial-time algorithms cannot do better than $(1+o(1))\frac{n\ln d}{d}$ (Coja-Oghlan & Efthymiou, 2015). Thus it is natural to measure the goodness of an algorithm by the ratio of the MIS size obtained to $\frac{n\ln d}{d}$. This ratio can faciliate comparison across different $n$ and $d$’s.
Figure 3 plots this ratio for several algorithms. Surprisingly for ReduMIS and OnlineMIS this ratio is consistently larger than $1.2$ and often larger than $1.3$ even for fairly large $n,d$. These results seem to bring the conjecture in doubt, but not disprove per se since the conjecture concerns asymptotically large $n$ and $d$. But our results could encourage further analysis and potential collaboration between researchers on theoretical and empirical aspects of MIS on random graphs.
Figure 3: Heatmap for ratios of MIS size to $\frac{n\ln d}{d}$ on ER graphs. We find that the ratio for ReduMIS and KaMIS is consistently larger than 1.2, suggesting that ReduMIS and KaMIS might surpass the conjectured upper bound (Coja-Oghlan & Efthymiou, 2015).
5 Related Works
Classical and heuristic methods for MIS
Classical methods for MIS range from simple greedy algorithms to advanced solvers like KaMIS which involves a number of heuristics. There are various existing heuristics for MIS, such as reduction techniques (Butenko et al., 2002; Xiao & Nagamochi, 2013; Akiba & Iwata, 2016), local search (Andrade et al., 2012), and evolutionary algorithms (Back & Khuri, 1994; Borisovsky & Zavolovskaya, 2003; Lamm et al., 2015). KaMIS (Lamm et al., 2017; Dahlum et al., 2016; Hespe et al., 2019) was developed based on these heuristics. In addition, MIS can be formulated into a binary integer programming problem (Nemhauser & Trotter Jr, 1975), which can be solved by the state-of-art integer programming solver Gurobi Gurobi Optimization, LLC (2024).
Machine learning for combinatorial optimization
In recent years, various ML-based algorithms have been developed for the MIS problem and most of them are based on graph neural networks (GNNs). Some of them using supervised learning (Li et al., 2018; Sun & Yang, 2023; Li et al., 2024) and requires labelling training data using classical solvers. Alternatively, those based on reinforcement learning (Khalil et al., 2017; Ahn et al., 2020; Qiu et al., 2022; Sanokowski et al., 2023) and other unsupervised learning objectives (Karalias & Loukas, 2020; Sun et al., 2022; Zhang et al., 2023; Sanokowski et al., 2024) do not require labeled training data. Zhang et al. (2023) (LTFT) uses GFlowNets (Bengio et al., 2021a) which is related to reinforcement learning. Notably, Ahn et al. (2020); Sanokowski et al. (2023); Zhang et al. (2023) model MIS problem as a Markov Decision Process (MDP) and generate the solution step-by-step (autoregressively). While Sanokowski et al. (2023) fixes the order of node updates, Ahn et al. (2020) (LwD) and Zhang et al. (2023) (LTFT) choose which node to update at each step so that they have a “natural” or “pseudo-natural” serialization and most suitable for our analysis in Section 4.1 and Section C.3. In addition, Sun & Yang (2023) (DIFUSCO) and Sanokowski et al. (2024) (DiffUCO) both utilizes diffusion model.
Most of the algorithms above also work on other types of graph CO problems, such as Maximum Cut, Maximum Vetex Cover, and Minimum Dominating Set. Some (Khalil et al., 2017; Qiu et al., 2022; Sun & Yang, 2023) also work for the Travelling Salesman Problem (TSP).
Other GPU-based solvers for CO
Recently, some non-learning GPU-based solvers for CO problems have been developed. Sun et al. (2023) developed a GPU-accelerated sampling based method which works on MIS, Max Cut, and TSP. Schuetz et al. (2022); Ichikawa (2023) uses GNNs conduct non-convex optimization for MIS without machine learning. Alkhouri et al. (2024) uses direct quadratic optimization without GNNs for MIS.
Benchmarks for MIS
Böther et al. (2022) provides a benchmark for several MIS algorithms including Gurobi (Gurobi Optimization, LLC, 2024), KaMIS (ReduMIS(Lamm et al., 2017), Intel-Treesearch (Li et al., 2018), DGL-Treesearch, and LwD Ahn et al. (2020). It is the only MIS benchmark we know about including recent AI-inspired method, though it only focuses its comparsion for learning-based tree search algorithms. It suggests that LwD is better than learning-based tree search algorithms. This aligns with our results where LwD performs the best among learning-based algorithms. This benchmark covers various types of random graphs and several real-world datasets, so it is a good reference benchmark for comparison over different types of graphs. Though unlike our benchmark, it does not provide comparison across various size and density for random graphs. Our benchmark fills this gap and provides a more detailed comparison. We also include many newer AI-inspired algorithms, and greedy algorithms which leads to detailed analysis like serialization. Angelini & Ricci-Tersenghi (2023) provides a comparison between Schuetz et al. (2022) and Deg-Greedy over regular graphs. The algorithm papers also report experiments comparing with previous algorithms, but those comparisons usually only focus on a few datasets and a few selected baselines.
6 Conclusion and Takeaways
Given the great interest in designing “general purpose AI reasoners”, it is interesting to check how well recent AI-based methods have fared in combinatorial optimization, a field with a long history of ingenious hand-designed algorithms. Our careful empirical comparisons of such AI-inspired methods with classical methods on MIS problem showed that none of the new methods outperform ReduMIS, the best CPU-based MIS solver, which builds up the independent set iteratively, sometimes backtracking (i.e. delete a vertex from the current set). As the graphs get larger or denser, the superiority of ReduMIS becomes more evident, whereas several AI-inspired algorithms drop to performing no better than trivial classical algorithms such as Deg-Greedy.
Further analysis shows that the fact that AI-inspired algorithms like DIFUSCO and PCQO select the independent set in one shot may be handicapping themselves by foregoing the benefits of local search. Methods like Deg-Greedy and LTFT select one node at a step and do not backtrack end up performing weakly. Interestingly, the best-performing learning-based algorithm LwD is doing something in the middle: selecting several nodes at a time, which may allow a more effective approximation to local search.
Using local search as post-processing improves the solutions of AI-inspired methods more than the CPU-based algorithms such as degree-based greedy. Despite that, with the local search step, these AI-inspired methods still perform worse than ReduMIS. This suggests that a more promising method to use AI and machine learning techniques for combinatorial optimization might be to teach the model the value of local search instead of simply use it as a post-processing.
Acknowledgement
YW, HZ, and SA acknowledge funding from DARPA. The authors would like to thank Alvaro Velasquez, Ismail Alkhouri, Kaifeng Lyu, Sadhika Malladi, and Pravesh Kothari for helpful discussions.
Appendix
Appendix A More Related Works
More on theoretical results for MIS
For random ER graphs with number of nodes $n$ and average degree $d$, the upper bound of MIS size is $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$(Coja-Oghlan & Efthymiou, 2015). Grimmett & McDiarmid (1975) proves that the simplest Ran-Greedy can achieve half-optimal at $\frac{n\ln d}{d}$ on random ER graphs. Despite that, there is no known existing polynomial algorithm which can reach MIS size of $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ for asymptotically large $n$ and $d$(Coja-Oghlan & Efthymiou, 2015). Coja-Oghlan & Efthymiou (2015) also suggests that the reason is likely independent sets of size larger than $(1+{\epsilon})\frac{n\ln d}{d}$ forms an intricately ragged landscape, where local algorithms will stuck. Gamarnik & Sudan (2014); Rahman & Virág (2017) proves that for local algorithms (which is defined to only use information from a constant neighborhood of a node to decide whether the node is in the independent set) are at most half optimal for independent set on random $d$-regular graphs. Gamarnik & Sudan (2014) suggests this is due to a property of the MIS problem, which they denote as the Overlap Gap Property (OGP) (Gamarnik, 2021). Gamarnik (2023) suggests that graph neural networks (GNNs) are also a type of local algorithms and thus being limited by OGP. While most AI-inspired MIS algorithms use GNN, the proof only applies to algorithms use GNNs as the only component to find solutions like (Schuetz et al., 2022), so it may not apply directly to more complicated algorithms like those tested in our paper. Yet, it may still suggests a reason why GNN-based algorithms (including most AI-inspired algorithms) cannot outperform classical heuristics like KaMIS.
In addition, Barbier et al. (2013) provides a conjectured tighter upper bound than $\frac{2n\ln d}{d}$ for $d$-regular graphs using the hard-core model in physics. Ding et al. (2016) proves a similar tighter upper bound for $d$-regular graphs. Wormald (2003) gives average-case performance for Deg-Greedy on $d$-regular graphs.
More on classical heuristics
Over the past few decades, significant progress has been made in tackling NP-hard combinatorial optimization (CO) problems by developing approximation algorithms and heuristic methods. Approximation algorithms provide provable guarantees on solution quality and have led to groundbreaking results for classical problems, such as the Maximum Independent Set (MIS), Traveling Salesperson Problem (TSP), and Maximum Cut (Boppana & Halldórsson, 1992; Laporte, 1992; Goemans & Williamson, 1995).
As we mentioned in Section 5, there are various existing heuristics for MIS. For exmaple, reduction techniques reduce the graph into smaller instances. Akiba & Iwata (2016) and Xiao & Nagamochi (2013) have shown a variety of reduction techniques which work well for MIS problem. Butenko et al. (2002); Bourgeois et al. (2012) uses reduction techniques to develop efficient exact algorithms for MIS. Local search improves an existing independent set by removing a small number of nodes and insert other eligible nodes. Andrade et al. (2012) gives an efficient local search algorithm for MIS and has been used as a subprocess for several MIS solvers. Evolutionary algorithms combine several existing solutions into a new solution. Examples include Back & Khuri (1994); Borisovsky & Zavolovskaya (2003); Lamm et al. (2015). KaMIS (Lamm et al., 2017; Dahlum et al., 2016) was developed based on many of these techniques above.
In addition, the MIS problem can also be relaxed into semi-definite programming (SDP), which leads to several approximation algorithms (Halperin, 2002; Bansal, 2014).
More on AI methods for combinatorial optimization
In addition to Section 5, we note that Sun et al. (2022); Sanokowski et al. (2023; 2024) use annealing techniques. Sun & Yang (2023) was developed based on Qiu et al. (2022), and Sanokowski et al. (2024) was based on Sanokowski et al. (2023). Sanokowski et al. (2024) can also be considered as an extension of Karalias & Loukas (2020).
Besides MIS, Maximum Cut, and TSP, people also considered to use AI-Inspired methods for other combinatorial optimization problems, including Vehicle Routing Problems (including TSP) (Kool et al., 2018; Chen & Tian, 2019; Delarue et al., 2020; Li et al., 2021; Zheng et al., 2021; Ye et al., 2024), Job Scheduling Problems (Lin et al., 2019; Baer et al., 2019; Zhang et al., 2020; Ye et al., 2024), Boolean Satisfiability (SAT) (Amizadeh et al., 2019; You et al., 2019; Kurin et al., 2020; Li & Si, 2022; Li et al., 2023), and Casual Discovery (Zheng et al., 2018; Zhu et al., 2020; Sanchez et al., 2022).
Appendix B Detailed Experiment Setup
B.1 Algorithm Pseudo-code for degree-based serialization
Please refer to Algorithm 1 for the pseudo-code for degree-based serialization, mentioned in Section 4.2.
Algorithm 1 Degree-based solution serialization
0:  The graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$, the independent set ${\mathcal{I}}$
1:  // Maintain the copy
2:  ${\mathcal{G}}^{\prime}\leftarrow{\mathcal{G}},{\mathcal{I}}^{\prime}\leftarrow{\mathcal{I}}$
3:  // Initialize empty ordered list
4:  ${\mathcal{L}}=[]$
5:  while ${\mathcal{I}}^{\prime}\neq\phi$ do
6:     pick $v\in{\mathcal{I}}^{\prime}$ with smallest degree in ${\mathcal{G}}^{\prime}$, break ties randomly
7:     ${\mathcal{L}}\leftarrow{\mathcal{L}}\cup\{v\}$
8:     Delete $\{v\}\cup\texttt{Neighbors}(v)$ from ${\mathcal{G}}^{\prime}$, delete $v$ from ${\mathcal{I}}^{\prime}$
9:  end while
10:  return  Ordered list ${\mathcal{L}}$
B.2 Datasets
For synthetic graphs, we test $8$ graphs for each parameter $(n,d)$ or $(n,m)$. We test on $100$ graphs for real-world datasets. For learning-based algorithms, we use $4000$ training graphs generated using the same parameter (in case of random graphs) or drawn $4000$ graphs from the same real-world dataset.
Erdős-Renyí (ER) graph
ER graphs (Erdös & Rényi, 1959) have 2 parameters: the number of nodes $n$ and the average degree $d$. For graphs with given $(n,d)$ We generate them by choosing $M=\frac{nd}{2}$ edges uniformly at random between $n$ nodes. This is the Erdős-Renyí’s $G(n,M)$ model.
There is also $G(n,p)$ model for ER graphs, which is also used widely. They behave similarly for many graph properties to $G(n,M)$ models when $M={n\choose 2}p$ and we expect the emipirical results for MIS problems will also be very similar.
For the main experiments, we use ER graphs with $n=\{100,300,1000,3000,10000\},d=\{10,30,100,300,1000,3000\}$ with $d<n$ as shown in Table 2. We also test on larger graphs with $n=\{30000,1\times 10^{5},3\times 10^{5},1\times 10^{6},3\times 10^{6}\}$ as shown in Table 7. Due to computational and time limits, we could only obtain results from classical CPU-based algorithms for these large graph, and only sparse graphs for very large $n\geq 1\times 10^{6}$.
Barabási–Albert (BA) graph
BA graphs (Albert & Barabási, 2002) also have 2 parameters: the number of nodes $n$ and the generation parameter $m$, with $n>m$. For a given $(n,m)$, the generation process initializes with $m$ nodes, and then add 1 node at each step. When adding a new node, $m$ neighbors of the new node are sampled from the existing nodes, with the probability of the current degree of the nodes. The average degree of BA graph with given $(n,m)$ can be computed as $d=2m-\frac{m}{n}-\frac{m^{2}}{n}\approx 2m$.
For the main experiments, similar to ER graphs, we also use $n=\{100,300,1000,3000,10000\}$. Since the average degree $d\approx 2m$, we use $m=\{5,15,50,150,500,1500\}$ with $2m<n$.
Real-world graphs
Since we need at least $4100$ graphs to train and test our algorithms and the graphs should not be too small, it is difficult to find such datasets. Fortunately, Morris et al. (2020) provides a website www.graphlearning.io which includes many graph datasets prepared by them or collected from other works. Although most of the datasets are still too small or having too small graphs, we are able to find 2 datasets: REDDIT-MULTI-5K and COLLA, both from Yanardag & Vishwanathan (2015). REDDIT-MULTI-5K has $4999$ graphs with average nodes $508.52$ and average edges $594.87$, so the graphs are generally very sparse. COLLA has $5000$ graphs with average nodes $74.49$ and $2457.78$, so they are denser but smaller graphs.
We were not able to find real-world datasets with enough size of more larger and denser graphs, which are generally more difficult for MIS algorithms. The dataset DIMACS used in Böther et al. (2022) contains such graphs but they only have $37$ graphs which is not enough for training.
B.3 Hardware configurations
The CPU we use is either Intel Xeon Processor E5-2680 v4 @ 2.40GHz or Intel Xeon Silver 4214 Processor @ 2.20GHz. The GPU we use is Nvidia Tesla A100 80GB when we refer to time limit or time cost. For small graphs when the GPU memory limit and time limit is not reached, we also use Nvidia RTX A6000 48GB and Nvidia RTX 2080Ti 11GB.
B.4 Classical CPU-based algorithms
Ran-Greedy and Deg-Greedy
Deg-Greedy (Degree-based Greedy) is as follows: Starting from an empty set. Select a node with the lowest degree of the graph (if there exists several nodes of the lowest degree, we pick one uniformly at random) and add it to the set. Then remove the node and all its neighbors from the graph.
Ran-Greedy (Random Greedy) is as follows: Starting from an empty set. Select a node from the graph uniformly at random and add it into the set. Then remove the node and all its neighbors from the graph. The only difference between it and Deg-Greedy is that the choice of node is completely random.
They are both non-backtracking algorithms.
We did not include results of Ran-Greedy in the main paper because its performance is significantly lower than all other algorithms. It can be considered as a baseline and has theoretical significance, since it has provable guarantee for random graphs (Grimmett & McDiarmid, 1975).
In order to match the best-of-20 sampling we used for the learning-based-algorithms, we also ran Deg-Greedy for $20$ times and report the best results in the main experiments ($n\leq 10000$).
We wrote the script in Python and ran it on a single CPU thread for each graph with a time limit of 24hrs, but it actually run less than 1hr on smaller graphs like $n\leq 3000$. We gave 32GB memory for graphs with $n\leq 10000$ and 64GB memory for larger graphs. Since there are much more efficient implementations like C++, the efficiency of the greedy algorithms is not very relevant.
KaMIS (OnlineMIS and ReduMIS)
KaMIS (Karlsruhe Maximum Independent Sets) (Lamm et al., 2017) (https://karlsruhemis.github.io/) is the state-of-art heuristic solver for MIS and has been used as a baseline in many previous works. It provides 2 algorithms for the MIS problem: ReduMIS (Lamm et al., 2017) and OnlineMIS (Dahlum et al., 2016).
We can provide a time-limit for both algorithm. OnlineMIS will use up the time given while ReduMIS will end on its own when it finds appropriate. In general, ReduMIS provides better results when given enough time, where OnlineMIS is faster to reach a solution of reasonable quality for large and dense graphs.
We ran both algorithm on a single CPU thread for each graph with a time limit of 24hrs, because our benchmark focus on performance instead of efficiency. For relatively small graphs ($n\leq 3000$), ReduMIS often require less than 1hr and at most $1.25$hrs, and OnlineMIS can also provide answer with the same quality when giving 1hr time limit. We gave 32GB memory for graphs with $n\leq 10000$ and 64GB memory for larger graphs.
B.5 GPU-accelerated non-learning algorithms
iSCO
iSCO (improved Sampling for Combinatorial Optimization) (Sun et al., 2023) is a GPU-accelerated sampling-based method. It does not require learning. According to (Sun et al., 2023), the main benefit is that it can process a large batch of graphs in parallel thus improve efficiency. While processing a small number of graphs like in our case (8 test graphs), it still requires significant time, often longer than ReduMIS.
We use the code from the codebase of DISCS (Goshvadi et al., 2024), which is a follow-up paper for iSCO, as iSCO did not provide the codebase. We use 1 80GB $A100$ GPU to run iSCO with all test graphs together. The time limit is 96hrs, and the actual time it requires is shorter. It fail to run graphs of size $(n=3000,1000)$, $(n=10000,d=100)$, and larger because they require larger than 80GB memory. The code does not support multi-GPU.
PCQO
PCQO (Parallelized Clique-Informed Quadratic Optimization) (Alkhouri et al., 2024) is a GPU-accelerated gradient-based optimization algorithm, which directly optimize on the quadratic loss function of the MIS problem. The loss function for a graph $G=(V,E)$ and solution vector ${\bm{x}}$ is
$f({\bm{x}})\coloneqq-\sum_{v\in V}{\bm{x}}_{v}+\gamma\sum_{(u,v)\in E}{\bm{x}}_{v}{\bm{x}}_{u}-\gamma^{\prime}\sum_{(u,v)\in E^{\prime}}{\bm{x}}_{v}{\bm{x}}_{u},$ (1)
where $E^{\prime}$ is the edge set of the complement graph $G^{\prime}$. $\gamma$ and $\gamma^{\prime}$ are hyperparameters and there are many other hyperparameters including learning rate, momentum, etc.
This algorithm is sensitive to hyperparameters and the default hyperparameters lead to bad solution quality for most of our dataset. Therefore, unlike in other algorithms where we use the default setting, we perform a hyperparameter tunining. We also did not find a good set of hyperparameters for all ER graphs or all BA graphs, so we do a grid search of hyperparameters for each dataset (i.e. for each $(n,d)$ pair in ER graphs and each $(n,m)$ pair in BA graphs).
Although Alkhouri et al. (2024) reported that using the third term in the loss function (i.e. nonzero $\gamma^{\prime}$) can lead to better answers comparing to only using 2-term loss function (set $\gamma^{\prime}=0$), we found that a nonzero $\gamma^{\prime}$ makes the algorithm more sensitive to hyperparameters and difficult to perform hyperparameter tuning for our benchmark with many datasets. Therefore, we only use $2$-term loss function by setting $\gamma^{\prime}=0$.
In the grid search for hyperparameters, we use learning rate $0.001$ (default), momentum $\{0.5,0.9\}$, $\gamma\in\{4,20,100,500,2500\}$, and report the results obtained from the best hyperparameters. However, for $n\geq 30000$, the grid search within this domain does not provide solution with reasonable size (worse than Ran-Greedy), so we do not report results for larger graphs. Despite that, there may exist a better set of hyperparamters which could make this algorithm perform well on larger graphs.
We use 1 80GB $A100$ GPU for all test graphs together with a time limit of 96hrs. It is able to run all experiments from $n\leq 10000$, and often requires shorter time compared to iSCO and KaMIS.
B.6 Learning-based algorithms
We use $4000$ training graphs for each datasets to train our models. Without otherwise noted, all the training are in-distribution, with respect to a single set of parameters (i.e. $(n,d)$ for ER graphs, $(n,m)$ for BA graphs) for synthetic graphs.
LwD
LwD (Learning what to Defer) (Ahn et al., 2020) is a reinforcement learning based algorithm which requires training data to learn the policy. It models the MIS problem as a Markov Decision Process (MDP), where in each step it selects some (possibly 00, $1$, or multiple) nodes to add into the independent set. It is a non-backtracking algorithm as the added nodes are never taken out. Böther et al. (2022) also included this algorithm in their benchmark.
We use the code from Böther et al. (2022) since it provides better functionality than the original codebase. We use the default setting provided by Böther et al. (2022) in their MIS Benchmark codebase (https://github.com/MaxiBoether/mis-benchmark-framework), but change the number of samples to $20$ (default is $10$) for test sampling, in order to match the best-of-20 sampling in our benchmark.
We use 1 80GB $A100$ GPU to train for each datasets and test with the same GPU. The training time limit is set to 96hrs. The default number of training steps (number of updates to the policy) is $20000$. Since LwD stores checkpoints throughout the process, we still report the test results based on the newest checkpoints for unfinished experiments if we have the checkpoints which reports meaningful results (better than half of the results reported by Deg-Greedy). Those results are indicated by * in the tables. The number of steps taken by those datasets with unfinished experiments is in Table 5.
Table 5: Number of Steps at termination for unfinished LwD experiments
Type of Graphs Parameters Number of Steps at termination ER $(n,d)$ $(1000,300)$ 1500 $(3000,100)$ 7200 BA $(n,m)$ $(1000,50)$ 18600 $(1000,150)$ 9600 $(3000,15)$ 9300 $(3000,50)$ 600
LTFT
LTFT (Let the Flows Tell) (Zhang et al., 2023) similar to LwD, also model the MIS problem as a MDP and non-backtracking. The difference is that it only choose $1$ node at each step, making it more similar to Deg-Greedy. The node chosen at each step is chosen by a GFlowNet (Bengio et al., 2021a), which is trained by in-distribution training data.
We use the default setting provide by Zhang et al. (2023) for training with $20$ epochs. By default setting, it has best-of-20 sampling and report the best solution found.
We use 1 80GB $A100$ GPU to train for each datasets and test with the same GPU. The training time limit is set to 96hrs. It completes training for all graphs with $n\leq 3000$, but larger graphs require larger GPU memory. The code does not support multi-GPU.
DIFUSCO
DIFUSCO (Diffusion Solvers for Combinatorial Optimization) (Sun & Yang, 2023) trains a diffusion model using supervised learning to produce a solution for the MIS. The diffusion model provides an entire solution so it is a one-shot algorithm.
The training data is $4000$ graphs for each dataset (1 set of parameter for synthetic graphs). All training is in-distribution. The training data is labelled by ReduMIS with time limit of 1hr. For graphs we used for training ($n\leq 3000$), ReduMIS gives the same performance compared to a time limit of 24hrs.
We use the default setting in (Sun & Yang, 2023) except that we use $50$ diffusion steps throughout training and testing, and $20$ samples for testing to be aligned with best-of-20 sampling in other methods. We train the model for $50$ epochs (default) for each dataset.
We use 1 80GB $A100$ GPU to train for each datasets and test with the same GPU. The training time limit is set to 96hrs. The code does not support multi-GPU. We report results where the training can be completed.
Although Sun & Yang (2023) suggested that DIFUSCO has some generalization ability. We found the performance degrade significantly for out-of-distribution trained models (specifically trained on smaller graphs with the same average degree but test on larger graphs), we did not report the results of larger graphs where the in-distribution training cannot finish.
DiffUCO
DiffUCO (Diffusion for Unsupervised Combinatorial Optimization) (Sanokowski et al., 2024) is also a diffusion model based algorithm but unlike DIFUSCO it uses unsupervised learning. The diffusion model is trained to sample the solution of low energy state. It also provides an entire solution and is also a one-shot algorithm.
We use the default setting in Sanokowski et al. (2024) for RB-large MIS task (in their Appendix C.5). During testing, we use conditional expectation with $20$ samples to align with best-of-20 sampling in other algorithms. The code supports multi-GPU. We use 4 80GB $A100$ GPU to train for each datasets with time limit 96hrs.
[…]
Table 6: Parameters of Test and Training Graphs for out-of-distribution testing in DiffUCO
Type of Graphs Parameters of Test Graphs Parameters of Training Graphs ER $(n,d)$ $(1000,300)$ $(1000,100)$ $(3000,10)$ $(1000,10)$ $(3000,30)$ $(3000,30)$ $(3000,100)$ $(1000,100)$ BA $(n,m)$ $(1000,150)$ $(1000,50)$ $(3000,5)$ $(1000,5)$ $(3000,15)$ $(1000,15)$ $(3000,50)$ $(1000,50)$
B.7 Local search
Local search is a method to improve a given independent set. It can be used as a post-processing technique, or be used as sub-procedures in more complicated algorithms like KaMIS. Andrade et al. (2012) provides an efficient local search algorithm. Part of it is to find $2$-improvement, which is the part used as sub-procedure in KaMIS (Lamm et al., 2017; Dahlum et al., 2016).
The local search algorithm for $2$-improvement for a given independent set $I$ is as follows. This algorithm process every vertex $x\in I$ in turn. First, it temporarily removes $x$ from $I$, creating a new set $S$. We call a vertex a free vertex of $S$ if there is no edge between it and any vertex in $S$. If $S$ has less than two free vertices, stop: there is no 2-improvement involving $x$. Otherwise, for each neighbor $v$ of $x$ that is a free vertex for $S$, insert $v$ into $S$ and check if the new set ($S^{\prime}$) has a free vertex $w$. If it does, inserting $w$ leads to a $2$-improvement; if it does not, remove $v$ from $S^{\prime}$ (thus restoring $S$) and process the next neighbor of $x$. If no improvement is found, reinsert $x$ into $S$ to turn it back to $I$. Every vertex is scanned $O(1)$ times in this algorithm so it can find a $2$-improvement (if there exists) in $O(m)$ time according to Andrade et al. (2012).
We implemented this algorithm in Python and use it as a post-processing for the solutions produced by the algorithm we test.
Appendix C More Experiment Results
In this section, we show more experiment results. Section C.1 shows more experiment results on much larger graphs, where the AI-inspired methods cannot handle. Section C.2 show the serialization results on more graphs. Section C.3 show a more detailed results between LwD and Deg-Greedy, which applies degree-based serialization as a subprocedure. Section C.4 shows the full results when adding local search as a post-processing procedure. Finally, Section C.5 discusses more results on the ratio of the solution found by different algorithms to $\frac{n\ln d}{d}$ on ER graphs, which relates to the theoretical conjecture mentioned in Section 4.4.
C.1 Larger graphs
Table 7 reports our results for large ER graphs not reported in Table 1. Within our computation limits as described in Appendix B, we can only obtain results for classical heuristic algorithms (Ran-Greedy, Deg-Greedy, OnlineMIS, ReduMIS).
Table 7: (Comparison of the performance of different algorithms on larger Erdős–Rényi (ER) graphs) Continuation of Table 1 on ER graphs with number of nodes larger than 100,000 for classical heuristic algorithms. Other algorithms are out of our computational limits for these large graphs.
Heuristics $n$ $d$ Ran-Greedy Deg-Greedy OnlineMIS ReduMIS 30000 10 7170.50 8951.75 9486.00 9505.38 30 3429.38 4464.12 4832.88 4914.88 100 1386.00 1817.25 1963.12 2012.62 300 570.50 738.00 794.75 815.50 1000 205.63 260.75 279.88 287.50 3000 106.00 106.12 100000 10 23990.00 29856.38 31613.38 31650.62 30 11444.38 14870.12 16128.88 16287.50 100 4615.25 6064.50 6564.00 6702.38 300 1886.62 2470.75 2661.50 2714.75 1000 83.50 941.38 959.75 300000 10 71954.38 89487.12 94622.88 94799.75 30 34318.88 44653.38 48234.38 48713.00 100 13850.12 18190.38 19676.50 20061.38 300 5688.00 7987.50 8141.25 1000 2831.38 3000 802.00 1000000 10 239749.12 312462.62 315630.88 30 114397.00 158625.29 161622.75 100 46161.12 64915.00 66539.75 300 26458.75 1000 6890.00 9473.00 3000000 10 719348.75 904613.12 942475.57 30 343479.38 459671.29 479938.75
C.2 Serialization
Figures 4 and 5 shows the percentage to choose the smallest possible degree node on different part of the serialization across various algorithms for all ER graphs and all BA graphs with nodes $n\leq 3000$, respectively. The serialization process is discussed in Section 4.2. Missing bars are algorithms which we do not get results due to computational limit, same as in Tables 1 and 2.
Figure 4: The percentage to choose the smallest possible degree node on different part of the serialization for all ER graphs It reinforces our observations in Section 4.2. In addition, we also observe that algorithms similar to Deg-Greedy (Deg-Greedy and LTFT) and good-performing algorithms (OnlineMIS, ReduMIS, iSCO, and LwD) have clearly different characteristics across various $(n,d)$, described in Section C.2.
Figure 5: The percentage to choose the smallest possible degree node on different part of the serialization for all BA graphs It reinforces our observations in Section 4.2. In addition, we also observe that algorithms similar to Deg-Greedy (Deg-Greedy and LTFT) and good-performing algorithms (OnlineMIS, ReduMIS, iSCO, and LwD) have clearly different characteristics across various $(n,m)$, described in Section C.2.
These results reinforced our observations in Section 4.2. First, the percentage for Deg-Greedy and LTFT are generally high. Deg-Greedy reaches $100\%$ for all parts in some graphs, which is the theoretically achievable percentage since Deg-Greedy actually picks the lowest degree node in the remaining graph and this sequence will give a serialization with percentage $100\%$ for all parts. In those cases, LTFT also have percentage close to $100\%$ for all $3$ parts.
Second, for those algorithms with good performance, namely OnlineMIS, ReduMIS, iSCO, and LwD, the bar plot shows similar characteristics. The percentage for the 1st third is generally low, while the 2nd third is high, and the 3rd third is higher and close to $100\%$. This characteristics are observed in most settings accross various parameters $(n,d)$/$(n,m)$ for both ER and BA graphs. The exception is only very sparse BA graphs.
Moreover, newly from these plots across various parameters, we also observe that given same $n$, Deg-Greedy and LTFT tend to have lower percentage for sparse graphs (smaller $d$ or $m$) and higher percentage for denser graphs (larger $d$ or $m$) across all $3$ parts. On the other hand, OnlineMIS, ReduMIS, iSCO, and LwD tend to have the percentage of 1st 1/3 decreases, while the percentage of the 2nd and 3rd 1/3 increases, when the density of graph increases ($d$ or $m$ increases for same $n$). This shows another qualitative difference between the algorithms similar to Deg-Greedy (Deg-Greedy and LTFT) and the good-performing algorithms (OnlineMIS, ReduMIS, iSCO, and LwD).
We also note that BA graphs $(n=300,m=50)$, $(n=1000,m=150)$, and $(n=3000,m=500)$ are outliers. Most algorithms have percentage close to $100\%$ for all parts. This is because these graphs are rather different from other BA graphs. They have an easily found large MIS, which is the $m$ nodes initially in the graph at the start of the BA generation process. (Albert & Barabási, 2002). From Table 2, we can see many algorithms can find these MIS and report a MIS size of $m$ for these graphs. This suggests that for this special type of BA graphs, our serialization analysis can observe different characteristics from other BA graphs.
C.3 Comparison between Deg-Greedy and LwD
LwD, similar to LTFT, is also a non-backtracking MDP based algorithm which picks nodes sequentially. The main difference is that instead of picking $1$ node at a step like Deg-Greedy and LTFT, it picks some nodes at a step, which can be 00 or $1$ or multiple nodes. In that case, it does not have a natural serialization like LTFT (discussed in Section 4.1). However, since it still have steps and we still know some nodes are chosen before others, we can perform a serialization within each step. We call this pseudo-natural serialization.
The procedure of our pseudo-natural serialization is as follows. Consider a step $t$ of LwD, let the independent set before the step be ${\mathcal{I}}_{t-1}$. Similar to Algorithm 1, we build a residual graph ${\mathcal{G}}^{\prime}$ which removes the nodes in ${\mathcal{I}}_{t-1}$ and their neighbors from ${\mathcal{G}}$. Then, LwD chose a set of nodes ${\mathcal{S}}_{t}$ to add to the independent set. We then perform the serialization for the set ${\mathcal{S}}_{t}$ (replace ${\mathcal{I}}^{\prime}$ by ${\mathcal{S}}$ in Algorithm 1) and get a ordered list ${\mathcal{L}}_{t}$. ${\mathcal{L}}_{t}$ is the ordered list for a step. We then concatenate all the ordered lists ${\mathcal{L}}_{t}$’s for all the steps in order to get a full ordered list ${\mathcal{L}}$. For this serialization, we do not repeat it as in Algorithm 1.
Similar to Section 4.1, we plot a heatmap Figure 6 across various parameters for ER and BA graphs on the average percentage of the nodes being the smallest degree node in the residual graph. In addition to that, similar to Section 4.2, we divide the ordered list ${\mathcal{L}}$ into $3$ equal parts to compute the average percentage of the nodes being the smallest degree node seperately.
Figure 6: Percentage of the smallest possible degree node in the pseudo-natural serialization of LwD, i.e., behaves similarly to degree-based greedy, for the 3 equal parts of the serialization, and the overal average From the overall heatmaps, we can see LwD is not very similar to Deg-Greedy like LTFT. From the heatmaps for different 1/3 parts, we can see the percentage increases from the 1st 1/3 to the 3rd 1/3 for all the datasets (different parameters of the synthetic graphs). This aligns with our “counterfactual” serialization results for LwD in Section 4.2, where we also observe the percentage increases clearly from 1st 1/3 to 3rd 1/3.
The heatmaps for overall percentages suggest that LwD is not very similar to Deg-Greedy like LTFT. From the heatmaps for different 1/3 parts, we can see the percentage increases from the 1st 1/3 to the 3rd 1/3 for all the datasets (different parameters of the synthetic graphs). This aligns with our “counterfactual” serialization results for LwD in Section 4.2, where we also observe the percentage increases clearly from 1st 1/3 to 3rd 1/3. This shows that our serialization method in Section 4.2 can reflect the pattern correctly.
The percentages here in the heatmap is smaller than the percentages for “counterfactual” serialization in the bar graphs in Section 4.2. This is likely due to the fact that in the “counterfactual” serialization we repeat the serialization process for $100$ times and report the highest percentages we get, while here we only do serialization once for each step.
C.4 Local search
Table 8: Adding local search as a post-processing procedure. This is the full graph for local search described in Section 4.3.
Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER Graphs 100,10 29.50 (0.25) 30.50 30.50 30.62 (0) 27.00 (2) 30.38 (0) 29.00 (0.38) 30.25 (0) 100,30 13.62 (0) 14.00 14.75 14.50 (0) 12.00 (0.50) 14.38 (0) 13.25 (0.13) 13.88 (0) 300,10 92.12 (0.50) 93.88 94.38 94.75 (0) 91.50 (0.88) 94.25 (0) 90.12 (1.50) 93.50 (0) 300,30 44.75 (0.25) 47.88 47.88 47.62 (0) 43.00 (1.25) 46.88 (0) 44.00 (0.75) 44.62 (0.74) 300,100 16.12 (0) 18.00 18.38 18.00 (0) 14.25 (0.87) 17.12 (0.12) 16.25 (0) 16.88 (0.26) 1000,10 305.38 (2.13) 314.75 316.13 315.62 (0) 308.12 (3) 311.25 (0) 300.00 (3.00) 306.38 (2.50) 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 148.50 (5.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 59.00 (3.38) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 1000,300 22.25 (0) 25.00 25.75 24.62 (0) 21.25 (1.25) 20.75 (1.63) 22.62 (0) 21.62 (0.74) 3000,10 913.62 (6.50) 947.25 954.25 950.88 (0) 927.00 (8.00) 935.38 (1.26) 900.62 (12.37) 918.00 (16) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 453.88 (11.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) 3000,100 185.62 (2) 194.38 201.50 200.38 (0) 180.25 (9.87) 191.50 (0.75) 185.62 (1.62) 182.25 (10.87) 3000,300 74.00 (0.50) 77.63 80.75 78.88 (0) 68.00 (4.50) 74.38 (0.50) 3000,1000 23.38 (0) 26.00 26.25 21.38 (1.50) 23.62 (0) BA Graphs 100,5 39.25 (0) 39.50 39.50 39.50 (0) 36.62 (2.74) 39.50 (0) 38.50 (0.38) 39.38 (0) 100,15 21.00 (0) 21.63 21.63 21.62 (0) 18.88 (0.50) 21.62 (0) 20.62 (0) 21.25 (0) 300,5 122.38 (0.26) 123.13 123.13 123.12 (0) 120.25 (2.63) 123.12 (0) 118.62 (3.00) 123.00 (0.12) 300,15 70.00 (0.75) 71.38 71.38 71.38 (0) 67.62 (2.00) 70.75 (0) 66.62 (1.87) 70.00 (0.50) 300,50 39.25 (1.75) 49.88 50.00 50.00 (0) 32.38 (5.13) 50.00 (0) 43.62 (1.87) 50.00 (0) 1000,5 413.38 (2) 417.13 417.13 417.12 (0) 412.62 (1.74) 416.00 (0) 400.25 (14.50) 417.12 (0) 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 235.50 (4.50) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 110.25 (2.37) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 1000,150 90.88 (7.63) 150.00 150.00 150.00 (0) 112.75 (3.13) 150.00 (0) 87.62 (5.37) 3000,5 1241.50 (4.75) 1257.00 1257.13 1255.62 (0) 1242.12 (3.37) 1248.50 (0.38) 1213.12 (35.87) 1254.75 (0.37) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 723.50 (12.88) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37) 3000,50 339.00 (3.62) 362.63 369.75 368.25 (0) 348.50 (9.62) 341.00 (4.12) 334.00 (10.12) 3000,150 142.88 (2.38) 160.25 165.75 164.00 (0) 150.50 (6.50) 146.50 (1.75) 3000,500 172.12 (9.12) 500.00 500.00 500.00 (0) 229.62 (5.74)
Table 8 shows the full results after incorporating $2$-improvement local search from the ARW local search algorithm (Andrade et al., 2012) as a post-processing step, which is discussed in Section 4.3.
C.5 More results on the ratio
In addition to what we show in Section 4.4, Figure 7 shows the ratio of MIS size to $\frac{n\ln d}{d}$ for ER graphs with number of nodes $n$ and average degree $d$ across more algorithms. We can see that ReduMIS, OnlineMIS, and iSCO has consitently high ratios more than $1.2$. Ran-Greedy stays around $1.0$ for all $(n,d)$. Other algorithms, including Deg-Greedy, all have higher ratios for sparser graphs, but lower ratios (close to $1$) for denser graphs.
Figure 7: Heatmap for ratios of MIS size to $nln(d)/d$ on ER graphs with number of nodes $n$ and average degree $d$. We can see that ReduMIS, OnlineMIS, and iSCO has consitently high ratios more than $1.2$. Ran-Greedy stays around $1.0$ for all $(n,d)$. Other algorithms, including Deg-Greedy, all have higher ratios for sparser graphs, but lower ratios (close to $1$) for denser graphs.

### v2

@@
-
+Time to Rethink AI for Combinatorial Optimization: Classical Algorithms Remain Tough to Match
 Abstract
-AI methods, such as generative models and reinforcement learning, have recently been applied to combinatorial optimization (CO) problems, especially NP-hard ones.
-This paper compares such GPU-based methods with classical CPU-based methods on Maximum Independent Set (MIS).
-Experiments on standard graph families show that AI-based algorithms fail to outperform and, in many cases, to match the solution quality of the state-of-art classical solver KaMIS running on a single CPU.
-Some GPU-based methods even perform similarly to the simplest heuristic, degree-based greedy.
-Even with post-processing techniques like local search, AI-based methods still perform worse than CPU-based solvers.
-We develop a new mode of analysis to reveal that non-backtracking AI methods, e.g.
-LTFT (which is based on GFlowNets), end up reasoning similarly to the simplest degree-based greedy approach, and thus worse than KaMIS.
-We also find that CPU-based algorithms, notably KaMIS, have strong performance on sparse random graphs, which appears to refute a well-known conjectured upper bound for efficient algorithms from Coja-Oghlan & Efthymiou (2015).
+This position paper argues that the machine learning community should fundamentally rethink how AI-inspired methods are developed and evaluated for combinatorial optimization (CO).
+We present comprehensive empirical benchmarks comparing various recent AI-inspired GPU-based methods with several classical CPU-based methods solvers on the Maximum Independent Set (MIS) problem.
+Strikingly, even on in-distribution random graphs, leading AI-inspired methods are consistently outperformed by state-of-art classical solver KaMIS, and some AI-inspired methods frequently fail to surpass even the simplest degree-based greedy heuristic.
+To better understand the source of these failures, we introduce a novel analysis, serialization, which reveals that non-backtracking AI methods, e.g.
+LTFT (which is based on GFlowNets), end up reasoning similarly to the simplest degree-based greedy, and thus worse than KaMIS.
+Our findings reveal three core issues: (1) Limited benchmarks and evaluation—AI-inspired methods are often tested only on small instances with very limited inference time, which covers up issues with scalability and resource usage; (2) Intrinsic hardness and learning limits—even under ideal, in-distribution conditions, learning-based approaches lag behind classical heuristics, highlighting inherent barriers that receive little attention; and (3) Insufficient use and understanding of classical heuristics—current learning frameworks often neglect to incorporate effective classical techniques.
+Although we use MIS as testbed, similar gaps and challenges have been reported in other combinatorial optimization problems, suggesting broader relevance for our recommendations.
+We propose that future research must address these issues by rigorous benchmarking, deepening theoretical and empirical understanding of learning limitations, and integrating classical heuristics into AI-inspired methods in a principled, evidence-driven manner.
 1 Introduction
-Combinatorial optimization (CO) lies at the core of numerous scientific and engineering studies, encompassing applications in network design, resource allocation, healthcare, and supply chain (Du & Pardalos, 1998; Hoffman, 2000; Zhong & Tang, 2021).
-Combinatorial optimization usually involves selecting an optimal solution from a discrete but often exponentially large set of candidates.
-Many are NP-hard (meaning that if $P\neq NP$ then there is no polynomial-time algorithm for solving them in the general cases (Papadimitriou & Steiglitz, 1998)).
-This makes it challenging to design algorithms with provable guarantees, but in practice solvers can find reasonable quality solutions (e.g., Gurobi (Gurobi Optimization, LLC, 2024)).
-Recent advances in artificial intelligence (AI) and GPU computing have motivated use of AI-inspired approaches, e.g., Graph Neural Networs (GNNs) and reinforcement learning, to learn problem-specific strategies for NP-hard optimization problems such as MIS (Li et al., 2018; Ahn et al., 2020) and TSP (Kool et al., 2018; Zhang et al., 2021).
-AI models can also be trained to predict search directions or refine heuristic rules (Li et al., 2018; d O Costa et al., 2020).
-These algorithms utilize advanced GPUs and often require shorter inference compared to CPU-based algorithms.
-Additionally, AI-inspired methods avoid the need to design heuristics for specific problems, allowing generalization to new instances and problems (Bengio et al., 2021b; Cappart et al., 2023).
-Despite the claimed benefits of AI-inspired methods, a few years ago Angelini & Ricci-Tersenghi (2023) showed that a Graph Neural Network based MIS algorithm (Schuetz et al., 2022) failed to surpass greedy algorithms.
-Böther et al.
-(2022) showed that AI-inspired approaches fail to provide superior search directions compared to traditional heuristics in tree search algorithms for MIS.
-Gamarnik (2023) suggests that GNN has theoretical limits which may become obstacles for GNN-based MIS algorithms.
-However, in recent years many new AI-inspired CO algorithms, utilizing a variety of techniques, such as diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024), GPU-accelerated sampling (Sun et al., 2023), and direct optimization (Alkhouri et al., 2024) have been developed and claimed to significant improve over the previous ones.
-Furthermore, GFlowNets Bengio et al.
-(2021a), which have been proposed as general-purpose tools for tasks like scientific discovery and reinforcement learning Bengio et al.
-(2023), has also been used to solve CO problems (Zhang et al., 2023).
-Since combinatorial optimization is an arena where humans have hand-designed algorithms for many decades, the following question is of great scientific interest:
-Do AI-inspired algorithms perform better than classical heuristics for
-combinatorial optimization?
-1.1 Our contributions
-We explore the question in the context of Maximum Independent Set (MIS) problem: given a graph, aiming to find the largest subset of nodes with no edges present between any node pair.
-The simplicity of the problem attracted design of many heuristics to tackle the problem (Andrade et al., 2012; Lamm et al., 2015).
-In recent works, MIS is also a main target in efforts that design AI-inspired approaches such as non-convex optimization (Schuetz et al., 2022; Alkhouri et al., 2024), reinforcement learning (Ahn et al., 2020; Zhang et al., 2023), and diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024).
-For classical heuristics, we test degree-based greedy (Deg-Greedy, pick a node with smallest degree at each step), and the state-of-the-art MIS solver KaMIS (Lamm et al., 2017; Dahlum et al., 2016; Hespe et al., 2019).
-For AI-inspired algorithms, we test the newest algorithms from each “category” according to the techniques they use, including sampling algorithm iSCO (Sun et al., 2023), non-convex optimization algorithm PCQO (Alkhouri et al., 2024), reinforcement-learning related algorithms LwD (Ahn et al., 2020) and LTFT (Zhang et al., 2023) (based on GFlowNets (Bengio et al., 2021a)) and diffusion models DIFUSCO (Sun & Yang, 2023) and DiffUCO (Sanokowski et al., 2024).
-Testing on different graph types with different sizes and densities leads to the following empirical finding (Section 3).
-Finding 1: Current AI-inspired algorithms for MIS still don’t outperform the best heuristic KaMIS, which runs on a single thread in a CPU, while AI-inspired methods often require significant computational resources.
-Finding 2: As the graph becomes larger or denser, KaMIS exhibits a notable superiority to AI-inspired algorithms.
-Finding 3: The simplest degree-based greedy algorithm (Deg-Greedy) serves as a very strong baseline.
-Some AI-inspired algorithms perform similarly to or worse than Deg-Greedy, especially for larger and denser graphs.
-Section 4 presents ablation studies to understand why some AI-inspired methods fail to improve over the simplest Deg-Greedy method.
-We introduce a new mode of analysis, serialization, that transforms the solution of any algorithm into a sequential list of choices leading to the final independent set.
-We compare sequential order with the one produced by Deg-Greedy (Sections 4.1 and 4.2).
-We find that the reinforcement learning related algorithm LTFT, based on GFlowNets, behaves similarly to Deg-Greedy.
-We also found several qualitative characteristics that appear to distinguish algorithms that perform significantly better than Deg-Greedy from those that perform similarly or even worse than Deg-Greedy.
-We also explore whether AI-inspired method can improve their solution quality via a post-processing step using local search (Section 4.3), but find that they still fail to outperform KaMIS after some improvements.
-In Section 4.4, we discuss an additional result that may be of interest for MIS experts: on random graphs, KaMIS has a level of performance on MIS that appears to contradict a well-known conjecture about polynomial-time algorithms Coja-Oghlan & Efthymiou (2015).
-2 Benchmarking MIS Algorithms
-We focus the experiment setup for benchmarking different algorithms for Maximum Independent Set problems (MIS).
+Combinatorial optimization (CO) lies at the core of numerous scientific and engineering studies, encompassing applications in network design, resource allocation, healthcare, and supply chain (du1998handbook, ; hoffman2000combinatorial, ; zhong2021preface, ).
+Most CO problems require selecting an optimal solution from a discrete but often exponentially large set of candidates and are typically NP-hard (papadimitriou1998combinatorial, ).
+While exact algorithms and advanced heuristics have long provided high-quality solutions in practice (e.g., Gurobi (gurobi, )), the last decade has seen a surge in machine learning (ML) and AI-inspired approaches to CO.
+These approaches—most notably, graph neural networks (GNNs), reinforcement learning (RL), and hybrid paradigms—promise to learn effective heuristics or even full solvers directly from data (li2018combinatorial, ; ahn2020learning, ; kool2018attention, ; zhang2021solving, ; bengio2021machine, ; cappart2023combinatorial, ).
+Early successes—such as pointer networks and RL-based methods for routing problems (kool2018attention, ), GNN-based frameworks for MIS and clique (khalil2017learning, ; li2018combinatorial, ), and hybrid algorithms combining classical solvers with learned advice (balcan2018learning, ; nair2020solving, )—fueled excitement that ML-driven solvers might soon rival or surpass specialized algorithms across a broad range of CO tasks.
+More recent AI-inspired algorithms, such as diffusion models (sun2023difusco, ; sanokowskidiffusion, ), GPU-accelerated sampling (sun2023revisiting, ), direct optimization (alkhouri2024dataless, ), and -general-purpose tools like GFlowNets (bengio2021flow, ; bengio2023gflownet, ; zhang2023let, ), have been claimed to continue these improvements.
+However, over the years, doubts have been raised about the practical gains of AI-inspired approaches.
+Empirical and theoretical studies raised concerns about their ability to generalize, to scale, and truly outperform strong classical baselines.
+A few years ago, (angelini2023modern, ) found that a GNN-based MIS algorithm (schuetz2022combinatorial, ) failed to outperform simple greedy baselines; (boether_dltreesearch_2022, ) showed that replacing heuristic components in tree search for MIS with GNN guidance offered little to no benefit, as performance largely stemmed from embedded classical reductions.
+Other works point to fundamental theoretical barriers for GNNs in CO (gamarnik2023barriers, ), and ML-based approaches for problems such as TSP have also struggled (joshi2021learning, ; xia2024position, ).
+These findings collectively suggested that it is difficult to replace classical heuristics with end-to-end learning and AI methods.
+Our paper evaluates more recent methods and our results suggest the same conclusion continues to hold.
+This paper explores similar issues using the latest AI methods.
+We choose in our comparative study on a single NP-complete problem, Maximum Independent Set (MIS) (“given an undirected graph find the largest subset of nodes with no edges present between any pair within the subset”).
+The reason for this choice are as follows.
+The MIS problem’s simple formulation has tempted researchers to try a range of solution strategies, from simple greedy heuristics to sophisticated heuristic solvers like KaMIS (lamm2017finding, ; hespe2019scalable, ).
+The MIS problem is interesting and nontrivial even on random graphs, and this allows a fair testing ground for AI-based methods, whose train/test paradigm is most directly suited to settings with a clean distribution on inputs.
+Unsurprisingly, AI-inspired methods of all sorts have been applied to MIS, including GNN (li2018combinatorial, ), RL (ahn2020learning, ; zhang2023let, ), and diffusion models (sun2023difusco, ; sanokowskidiffusion, ).
+We evaluate (Section 3) a broad suite of AI-inspired and classical MIS solvers on a variety of graph types, with an emphasis on in-distribution random graphs.
+The results (Section 4) are revealing: even on test graphs drawn from the same distribution as the ML models’ training data, the state-of-the-art classical algorithm KaMIS, which uses far less computational resources while finding larger independent sets.
+Even more surprisingly, some AI-inspired methods frequently fail to outperform even the simplest degree-based greedy heuristic.
+These findings demonstrate that current AI-inspired methods for MIS have yet to replicate, let alone surpass, the effectiveness of well-established heuristics -—sometimes not even the simplest one.
+The ML and AI community should rethink its approach to combinatorial optimization, focusing on the following issues (details in Section 6):
+(1) Limited benchmarks and unrealistic evaluation: Current research often evaluates AI-inspired methods on small, quick-inference instances, failing to capture scalability and resource usage.
+Future work should establish rigorous, realistic benchmarks and report computational costs.
+(2) Intrinsic hardness and learning limits: Even with well-defined distribution and entirely in-distribution training data, learning-based methods lag behind classical heuristics, indicating a need for deeper theoretical and empirical analysis of the limitations of learning algorithms.
+(3) Insufficient use and understanding of classical heuristics: Progress will require principled integration of effective classical heuristics into AI-inspired frameworks, informed by careful analysis.
+Although our testbed is MIS, we believe these positions broadly apply to other CO problems.
+Similar performance gaps and benchmarking issues have been found in TSP (joshi2021learning, ; sun2023difusco, ; xia2024position, ), and recent theoretical work highlights barriers for GNN-based methods across CO tasks (gamarnik2023barriers, ).
+To further support our position that AI-based methods are not providing novel insight into MIS problem, Section 5 presents ablation studies to understand why some AI-inspired methods fail to surpass the simplest Deg-Greedy method.
+We introduce a new mode of analysis, serialization, which reveals structural similarities between certain AI-driven methods and basic greedy heuristics (Sections 5.1 and 5.2).
+We also explore whether AI-inspired methods can improve their solutions through local search post-processing (Section 5.3), but find limited improvements.
+Our findings underscore both the intrinsic difficulty of MIS for learning-based approaches and the enduring value of classical heuristics.
+We conclude by outlining issues and future directions in Section 6, and also provide alternative views in Section 7.
+2 Related Works
+Classical and heuristic methods for MIS:
+Classical methods for MIS range from simple greedy algorithms to advanced solvers like KaMIS which involves a number of heuristics.
+There are various existing heuristics for MIS, such as reduction techniques (butenko2002finding, ; xiao2013confining, ; akiba2016branch, ), local search (andrade2012fast, ), and evolutionary algorithms (back1994evolutionary, ; borisovsky2003experimental, ; lamm2015graph, ).
+KaMIS (lamm2017finding, ; dahlum2016accelerating, ; hespe2019scalable, ) was developed based on these heuristics.
+In addition, MIS can be formulated into a binary integer programming problem (nemhauser1975vertex, ), which can be solved by the state-of-art integer programming solver Gurobi gurobi .
+Machine learning for combinatorial optimization:
+In recent years, various ML-based algorithms have been developed for the MIS problem.
+Most of them are based on graph neural networks (GNNs).
+Some use supervised learning (li2018combinatorial, ; sun2023difusco, ; li2024distribution, ) and requires labelling training data using classical solvers.
+Alternatively, those based on reinforcement learning (khalil2017learning, ; ahn2020learning, ; qiu2022dimes, ; sanokowski2023variational, ) and other unsupervised learning objectives (karalias2020erdos, ; sun2022annealed, ; zhang2023let, ; sanokowskidiffusion, ) do not require labeled training data.
+(zhang2023let, ) (LTFT) uses GFlowNets (bengio2021flow, ) which is related to reinforcement learning.
+Notably, (ahn2020learning, ; sanokowski2023variational, ; zhang2023let, ) model MIS problem as a Markov Decision Process (MDP) and generate the solution step-by-step (autoregressively).
+While (sanokowski2023variational, ) fixes the order of node updates, (ahn2020learning, ) (LwD) and (zhang2023let, ) (LTFT) choose which node to update at each step so that they have a “natural" or “pseudo-natural" serialization and most suitable for our analysis in Section 5.1 and Section C.5.
+(sun2023difusco, ) (DIFUSCO) and (sanokowskidiffusion, ) (DiffUCO) both use diffusion model.
+Most of the algorithms above also work on other types of graph CO problems, such as Maximum Cut, Maximum Vetex Cover, and Minimum Dominating Set.
+Some (khalil2017learning, ; qiu2022dimes, ; sun2023difusco, ) also work for the Travelling Salesman Problem (TSP).
+Other GPU-based solvers for CO:
+Recently, some non-learning GPU-based solvers for CO problems have been developed.
+(sun2023revisiting, ) developed a GPU-accelerated sampling based method which works on MIS, Max Cut, and TSP.
+(schuetz2022combinatorial, ; ichikawa2023controlling, ) uses GNNs conduct non-convex optimization for MIS without machine learning.
+(alkhouri2024dataless, ) uses direct quadratic optimization without GNNs for MIS.
+Benchmarks for MIS
+(boether_dltreesearch_2022, ) provides a benchmark for several MIS algorithms including Gurobi (gurobi, ), KaMIS (ReduMIS(lamm2017finding, ), Intel-Treesearch (li2018combinatorial, ), DGL-Treesearch, and LwD ahn2020learning .
+It is the only MIS benchmark we know about including recent AI-inspired method, though it only focuses its comparsion for learning-based tree search algorithms.
+This benchmark covers various types of random graphs and several real-world datasets, so it is a good reference benchmark for comparison over different types of graphs.
+Though unlike our benchmark, it does not provide comparison across various size and density for random graphs.
+Our benchmark fills this gap and provides a more detailed comparison.
+We also include many newer AI-inspired algorithms, and greedy algorithms which leads to detailed analysis like serialization.
+(angelini2023modern, ) provides a comparison between (schuetz2022combinatorial, ) and Deg-Greedy over regular graphs.
+The algorithm papers also report experiments comparing with previous algorithms, but those comparisons usually only focus on a few datasets and a few selected baselines.
+3 Benchmarking AI-inspired Methods for Maximum Independent Set
 3.1 Maximum Independent Set (MIS) problem
@@
 We classify the algorithms we test as: (1) classical heuristics, which includes Deg-Greedy and KaMIS (OnlineMIS and ReduMIS); (2) GPU-accelerated non-learning algorithms, which includes iSCO and PCQO; and (3) learning-based algorithms, which includes LwD, LTFT, DIFUSCO, and DiffUCO.
-Deg-Greedy
-(Degree-based greedy) Simplest heuristic: always picks a node with the smallest degree in the current graph, add to the current independent set, and delete that node and all of its neighbors from the graph.
+Deg-Greedy (Degree-based greedy) Simplest heuristic: always picks a node with the smallest degree in the current graph, add to the current independent set, and delete that node and all of its neighbors from the graph.
 Most papers on AI-inspired methods do not compare with this baseline.
-OnlineMIS and ReduMIS
-are two variants of the MIS solver KaMIS, mainly consists of three alternating steps: greedy, local search, and graph reductions.
-OnlineMIS (Dahlum et al., 2016) only applies a simple reduction after local search, while ReduMIS Lamm et al.
-(2017) applies many graph reduction techniques.
-iSCO
-(Sun et al., 2023) is a GPU-accelerated sampling-based method, incorporating gradient-based discrete MCMC and simulated annealing.
+OnlineMIS and ReduMIS are two variants of the MIS solver KaMIS, mainly consists of three alternating steps: greedy, local search, and graph reductions.
+OnlineMIS (dahlum2016accelerating, ) only applies a simple reduction after local search, while ReduMIS lamm2017finding applies many graph reduction techniques.
+iSCO (sun2023revisiting, ) is a GPU-accelerated sampling-based method, incorporating gradient-based discrete MCMC and simulated annealing.
 The MCMC is designed based on the Metropolis-Hasting algorithm, which if given enough time (exponential), can get the optimal solution.
-PCQO
-(Alkhouri et al., 2024) directly optimizes the quadratic loss function of the MIS using gradient descent.
+PCQO (alkhouri2024dataless, ) directly optimizes the quadratic loss function of the MIS using gradient descent.
 It is sensitive to optimization hyperparameters, so hyperparameter search is required for achieving good results.
-LwD
-(Ahn et al., 2020) is a reinforcement learning based algorithm which models the problem as a Markov Decision Process (MDP) and requires a dataset (without solutions) to train the policy.
+Extensive hyperparameter search may give better results than in our experiments.
+LwD (ahn2020learning, ) is a reinforcement learning based algorithm which models the problem as a Markov Decision Process (MDP) and requires a dataset (without solutions) to train the policy.
 In each step, several nodes are added to the independent set and are never deleted.
 We call it a non-backtracking algorithm.
-Ahn et al.
-(2020) reported that it outperforms KaMIS on very large but very sparse random graphs, which we do not include in our benchmark.
-LTFT
-(Zhang et al., 2023) is also a non-backtracking MDP-based algorithm similar to LwD, but it only selects one node at a time, which is decided by GFlowNets (Bengio et al., 2021a).
+LTFT (zhang2023let, ) is also a non-backtracking MDP-based algorithm similar to LwD, but it only selects one node at a time, which is decided by GFlowNets (bengio2021flow, ).
 Thus, it has a very similar procedure to Deg-Greedy.
 The algorithm requires a dataset (without solutions) to train the neural network.
-DIFUSCO and DiffUCO
-DIFUSCO (Sun & Yang, 2023) is an end-to-end one-shot MIS solver using diffusion models and requires a supervised training dataset with solutions.
-DiffUCO (Sanokowski et al., 2024) also uses diffusion model but with unsupervised learning and annealing techniques.
+DIFUSCO and DiffUCO DIFUSCO (sun2023difusco, ) is an end-to-end one-shot MIS solver using diffusion models and requires a supervised training dataset with solutions.
+DiffUCO (sanokowskidiffusion, ) also uses diffusion model but with unsupervised learning and annealing techniques.
 It requires a training dataset without solutions.
@@
 AI-inspired methos PCQO, DIFUSCO, and DiffUCO are one-shot algorithms, since they work like end-to-end MIS solvers and directly return the full set.
-[…]
-CPU-based GPU-acc Learning-based $n$ $d$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 10 29.25 30.50 30.50 30.62 25.00 30.38 28.62 30.25 30.02 30 13.63 14.00 14.75 14.50 11.50 14.38 13.12 13.88 13.92 300 10 77.50 93.88 94.38 94.75 90.62 94.25 88.62 93.50 91.84 30 44.50 47.88 47.88 47.62 41.75 46.88 43.25 43.88 45.04 100 16.13 18.00 18.38 18.00 13.38 17.00 16.25 16.62 16.93 1000 10 303.25 314.75 316.13 315.62 305.12 311.25 297.00 303.88 311.67 30 151.00 158.88 163.75 163.50 143.00 158.38 150.00 143.75 152.55† 100 60.63 64.75 66.63 66.50 55.62 63.88 60.88 55.38 63.55† 300 22.25 25.00 25.75 24.62 20.00 19.12* 22.62 20.88 18.36† 3000 10 907.13 947.25 954.25 950.88 919.00 934.12 888.25 902.00 935.28† 30 451.88 480.88 493.13 491.62 442.38 473.25 449.00 413.38 459.57† 100 183.63 194.38 201.50 200.38 170.38 190.75* 184.00 171.38 194.07† 300 73.50 77.63 80.75 78.88 63.50 73.88 1000 23.38 26.00 26.25 19.88 23.62 10000 10 2999.88 3161.88 3173.62 3149.92 2685.62 30 1498.00 1607.50 1639.88 1625.47 1334.75 100 613.75 650.00 670.88 573.63 300 249.00 258.38 272.25 239.13 1000 87.13 91.50 94.25 80.88 3000 29.63 33.00 33.25 19.00
-2.3 Graph types
+3.3 Datasets and experiment details
+We generated and experimented on 2 random graph types: Erdős-Renyí (ER) graphs (erdos59a, ) and Barabási–Albert (BA) graphs (albert2002statistical, ), across various sizes and densitites.
+We also experimented on existing datasets: synthetic RB graphs (xu2000exact, ) from (zhang2023let, ), and real-world datasets form (yanardag2015deep, ).
+Due to space limit, we only present ER graphs in main text.
+Descriptions of other graphs are in Section B.2.
 Erdős-Renyí (ER) graph
-(Erdös & Rényi, 1959) are random graphs where edges are connected uniformly at random (with a given probability or a fixed number of edges).
-We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing the number of edges at $\frac{nd}{2}$.
-Previous benchmark (Böther et al., 2022) and algorithms (Ahn et al., 2020; Sun & Yang, 2023; Zhang et al., 2023; Alkhouri et al., 2024) used it as test graphs for MIS, though without varying parameters as we did.
-Barabási–Albert (BA) graph
-(Albert & Barabási, 2002) are random graphs generated by a probabilistic growth process, mimicking real-world networks such as Internet, citation networks, and social networks (Albert & Barabási, 2002; Radicchi et al., 2011).
-For BA graphs, we vary $2$ parameters: number of nodes $n$ and parameter $m$ (not number of edges).
-The average degree of BA graphs can be approximated as $2m$.
-Real-world graphs
-We pick REDDIT-MULTI-5K and COLLAB (Yanardag & Vishwanathan, 2015) from TUDataset website (Morris et al., 2020), since they have enough graphs for training and graph sizes not too small.
-REDDIT-MULTI-5K has $508.52$ average nodes and $594.87$ average edges.
-They are mostly very sparse graphs.
-COLLAB has $74.49$ average nodes and 2457.78 average edges.
-They are mostly small but dense graphs.
-2.4 More experiment details
+(erdos59a, ) are random graphs where edges are connected uniformly at random (with a fixed probability or number of edges).
+We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing number of edges at $\frac{nd}{2}$.
+Previous works (boether_dltreesearch_2022, ; ahn2020learning, ; sun2023difusco, ; zhang2023let, ; alkhouri2024dataless, ) used it as test graphs for MIS, though without varying parameters as we did.
 For synthetic graphs, we test $8$ graphs for each parameter $(n,d)$ or $(n,m)$.
@@
 We set $24$-hr time limit for KaMIS (OnlineMIS and ReduMIS) since it runs on a single CPU thread with 32GB memory, and our benchmark focuses on performance on solution quality.
-Note that AI-based methods run well only on relatively small graphs, and ReduMIS runs in less than one hour on small graphs $(n\leq 3000)$.
+ReduMIS runs in less than one hour on small graphs $(n\leq 3000)$.
 Given 24 hours ReduMIS can handle much larger graphs (up to $n\approx 1e6$).
 For iSCO and learning-based algorithms, we report results within our computational limit (generally a single 80GB A100 for $96$hrs, details in Appendix B.) We test PCQO for $n\leq 10000$ because the performance degrades quickly for large graphs using default hyperparameter search domain.
-3 Experiment Results and Main Findings
+4 Performance Gap: Classical Methods Outperform AI-inspired Methods
 In this section, we present our main experiment results.
-The performance of different algorithms on Erdős-Renyí (ER) graphs, Barabási–Albert (BA) graphs, and real-world graphs are shown in Tables 1, 2 and 3 respectively.
+The results for various algorithms on Erdős-Renyí (ER) graphs are in Table 1.
+Due to space limit, the results for other graphs are in appendix: Barabási–Albert (BA) graphs in Table 5, RB graphs (xu2000exact, ) in Table 6, and real-world graphs in Table 7.
+[…]
+CPU-based GPU-acc Learning-based $n$ $d$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 10 29.25 30.50 30.50 30.62 30.63 30.38 28.62 30.25 30.02 30 13.63 14.00 14.75 14.50 14.50 14.38 13.12 13.88 13.92 300 10 77.50 93.88 94.38 94.75 94.63 94.25 88.62 93.50 91.84 30 44.50 47.88 47.88 47.62 47.63 46.88 43.25 43.88 45.04 100 16.13 18.00 18.38 18.00 18.00 17.00 16.25 16.62 16.93 1000 10 303.25 314.75 316.13 315.62 310.00 311.25 297.00 303.88 311.67 30 151.00 158.88 163.75 163.50 158.63 158.38 150.00 143.75 152.55† 100 60.63 64.75 66.63 66.50 60.13 63.88 60.88 55.38 63.55† 300 22.25 25.00 25.75 24.62 23.25 19.12* 22.62 20.88 18.36† 3000 10 907.13 947.25 954.25 950.88 923.25 934.12 888.25 902.00 935.28† 30 451.88 480.88 493.13 491.62 464.25 473.25 449.00 413.38 459.57† 100 183.63 194.38 201.50 200.38 185.63 190.75* 184.00 171.38 194.07† 300 73.50 77.63 80.75 78.88 69.25 73.88 1000 23.38 26.00 26.25 23.00 23.62 10000 10 2999.88 3161.88 3173.62 3149.92 2569.75 30 1498.00 1607.50 1639.88 1625.47 1331.63 100 613.75 650.00 670.88 543.13 300 249.00 258.38 272.25 218.63 1000 87.13 91.50 94.25 77.50 3000 29.63 33.00 33.25 26.63
 AI-inspired algorithms don’t outperform ReduMIS.
-Our first main finding is that, current AI-inspired algorithms do not outperform the best classical heuristics ReduMIS in terms of performance.
-As shown in Tables 1, 2 and 3, ReduMIS consistently achieves superior results compared to all other methods, with the exception of iSCO sometimes perform similarly, on ER, BA, and real-world graphs.
-Additionally, although learning-based algorithms are claimed to be more efficient, they require significant training time and GPU memory, which is over our resource constraint for graphs with more than 3000 nodes, while ReduMIS can handle graphs with up to $1\times 10^{6}$ nodes (See Table 7).
+Our first main finding is that, current AI-inspired algorithms do not outperform the best classical heuristic ReduMIS in terms of performance.
+As shown in Tables 1, 5 and 7, ReduMIS consistently achieves superior results compared to all other methods, with the exception of iSCO sometimes perform similarly.
+Although learning-based algorithms are claimed to be more efficient, they require significant training time and GPU memory, which is over our resource constraint for graphs with more than 3000 nodes, while ReduMIS can handle graphs with up to $1\times 10^{6}$ nodes (See Table 8).
 Although iSCO performs close to ReduMIS, it requires significant GPU memory and we are unable to get results for dense graphs with 10000 nodes.
@@
 In summary, our experiment results show that current AI-inspired algorithms still don’t outperform the best classical heuristics for the MIS problem.
-[…]
-Heuristics GPU-acc Learning-based $n$ $m$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 5 39.50 39.50 39.50 39.50 33.88 39.50 38.12 39.38 39.25 15 21.00 21.63 21.63 21.62 18.38 21.62 20.62 21.25 21.25 300 5 121.88 123.13 123.13 123.12 117.62 123.12 115.62 122.88 123.00 15 69.38 71.38 71.38 71.38 65.62 70.75 64.75 69.50 69.25 50 38.00 49.88 50.00 50.00 27.25 50.00 41.75 50.00 50.00 1000 5 412.63 417.13 417.13 417.12 410.88 416.00 385.75 417.12 416.00 15 232.75 245.00 246.38 246.25 231.00 243.12 224.25 236.38 240.75 50 107.50 115.75 116.88 116.75 107.88 113.00* 106.38 105.25 56.38† 150 81.50 150.00 150.00 150.00 109.62 150.00* 82.25 135.00† 3000 5 1236.63 1257.00 1257.13 1255.62 1238.75 1248.12 1177.25 1254.38 1252.63† 15 709.50 749.63 754.50 752.00 710.62 727.75* 661.88 726.38 737.38† 50 335.38 362.63 369.75 368.25 338.88 336.88* 323.88 142.25† 150 147.25 160.25 165.75 164.00 144.00 144.75 500 141.00 500.00 500.00 500.00 223.88 10000 5 4128.75 4206.00 4205.38 4190.47 4161.88 15 2377.88 2525.00 2534.00 2517.07 2428.13 50 1122.13 1228.88 1251.13 1241.98 1184.63 150 511.25 555.75 580.13 573.37 537.63 500 192.50 209.13 217.50 198.75 1500 67.25 1500.00 1500.00 1500.00
-Table 3: (Comparison of the performance of different algorithms on real-world graphs) Table 1 on real-world datasets.
-In general, graphs in REDDIT-MULTI-5K are very sparse, while COLLAB are dense but small.
-The numbers within ±$\pm 1\%$ of the best are highlighted.
-Dataset Heuristics GPU-acc Learning-based Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO REDDIT-MULTI-5K 350.73 350.73 350.66 350.73 344.47 350.73 343.35 350.72 350.69 COLLAB 8.68 8.70 8.70 8.70 8.57 8.69 8.70 8.75 8.70
 The performance gap between ReduMIS and AI-based methods widens with larger and denser graphs.
@@
 On ER graphs when $n=100,300$ and average degree $d=10$, LwD has results within $1\%$ gap from ReduMIS.
-On BA graphs for $n=300,1000$ and $m=5$, LwD, DIFUSCO, and DiffUCO has less than $0.5\%$ gap from ReduMIS.
-Across real-world datasets all tested algorithms produce independent sets of similar sizes on average.
 However, as the graph becomes larger or denser, the performance gap between ReduMIS and AI-based algorithms enlarges.
@@
 This gap widens further for $n=3000$, where AI-based algorithms perform significantly worse than ReduMIS and sometimes fail to outperform simple heuristic Deg-Greedy.
-A similar trend has also been observed on BA graphs.
 Classical solvers have no difficulty handling graphs with a million edges, but learning-based implementations struggle to scale up to that size.
@@
 Another key finding is that the simplest degree-based greedy (Deg-Greedy) serves as a remarkably strong baseline.
-As shown in Tables 1, 2 and 3, leveraging neural networks for node selection, LTFT often perform comparably to Deg-Greedy, particularly on larger or denser graphs.
+As shown in Table 1, leveraging neural networks for node selection, LTFT often perform comparably to Deg-Greedy, particularly on larger or denser graphs.
 For example on ER graphs when $n=1000$ or $3000$, LTFT gives performance within 2% of Deg-Greedy (Table 1).
-An exception is for BA graphs where the graph is very dense ($n=3000,m=500$) and clear patterns emerge in the graph, LTFT surpasses Deg-Greedy.
-Additionally, DIFUSCO and PCQO fail to outperform Deg-Greedy on larger or less sparse graphs, such as $n=3000$, $d>30$.
-4 AI-Inspired vs Deg-Greedy
+Additionally, DIFUSCO and PCQO fail to outperform Deg-Greedy on larger or denser graphs, such as $n=1000$ with $d\geq 100$ and $n=3000$ for DIFUSCO, $n=3000$ with $d\geq 300$and $n=10000$ for PCQO.
+Similar trends on other types of graphs
+We also tested on Barabási–Albert (BA) graphs (Table 5), RB graphs (Table 6), and real-world graphs (Table 7).
+For BA graphs and RB graphs, ReduMIS also outperform all other methods across various setting.
+For real-world graphs, most methods including ReduMIS perform close to the best, likely because these datasets are either small or sparse and thus easy for MIS problem.
+On BA and RB graphs, we also observe that the performance gap between ReduMIS and AI-inspired methods increase as the graphs become larger and denser.
+Deg-Greedy is also a strong baseline for all these graph types.
+Notably, for large RB graphs, all learning-based methods perform worse than Deg-Greedy.
+Since RB graphs are considered difficult cases for MIS (zhang2023let, ), these results may suggest learning-based methods are even weaker for graph types with higher "intrinsic hardness" than ER graphs.
+5 Deconstructing the Shortfall: Comparison with Deg-Greedy
 Our results show that the state-of-the-art AI-inspired algorithms for MIS still do not outperform the best heuristic ReduMIS.
 The surprising finding was that they also often do not outperform the simplest classical heuristic, Deg-Greedy, especially on large and dense graphs.
-In this section, we delve deeper into this comparison (Sections 4.1 and 4.2).
-Furthermore, we explore the impact of augmenting various algorithms with a local search as a post hoc step to enhance solution quality (Section 4.3).
+In this section, we delve deeper into this comparison (Sections 5.1 and 5.2).
+We also explore the impact of augmenting various algorithms with a local search as post processing to enhance solution quality (Section 5.3).
 5.1 Comparison between Deg-Greedy and LTFT
-Figure 1: Percentage of rounds when LTFT selects the node with smallest possible degree, i.e., behaves similarly to degree-based greedy.
+(a) Percentage of rounds when LTFT selects the node with smallest possible degree, i.e., behaves similarly to degree-based greedy.
 On ER graphs when the graph is dense (closer to the bottom left corner), the percentage of rounds LTFT selects the node with smallest possible degree is higher, i.e., behaves more similarly to degree-based greedy.
-On BA graphs, the percentage to choose the smallest possible node is generally lower, but on denser BA graphs LTFT also behaves more like degree-based greedy.
+(b) The percentage to choose the smallest possible degree node on different part of the (degree-based) serialization.
+We find that the best algorithms (OnlineMIS, ReduMIS, iSCO) share a similar characteristic pattern, that they have high consistency with degree-based greedy on second and third part of the serialization, while on the first part there is a low consistency.
+PCQO and LwD, which performs a bit worse, also share this pattern but a bit higher for the first part, and lower for second and third parts.
+By contrast, LTFT and DIFUSCO, which perform significantly worse, either has very high consistency for all parts(LTFT) or has low consistency for all parts (DIFUSCO).
+Figure 1: Analysis on ER graphs in (a) Section 5.1 and (b) Section 5.2
 Deg-Greedy sequentially picks nodes for the independent set.
@@
 As in Section 3.2 we call it a non-backtracking algorithm.
-LTFT is also a non-backtracking algorithm and it often perform similarly to Deg-Greedy in Tables 1 and 2.
-It uses a trained policy network GFlowNets (Bengio et al., 2021a) to pick a node for the independent set at each step.
+LTFT is also a non-backtracking algorithm and it often perform similarly to Deg-Greedy in Tables 1 and 5.
+It uses a trained policy network GFlowNets (bengio2021flow, ) to pick a node for the independent set at each step.
 Thus, we can naturally compare LTFT with Deg-Greedy by investigating how close this trained policy compares to the naive policy in Deg-Greedy.
-The results are shown in Figure 1.
+The results are shown in Figure 1(a).
 Overall, we observe that LTFT frequently selects nodes with very small degrees.
 On ER graphs where the average degree is at least 30, LTFT picks the node with the smallest degree over 85% of rounds except for sparse graphs ($d=10$ and $(n,d)=(3000,30)$).
-On BA graphs, while the percentage is lower, it still exceeds 75% on large and sufficiently dense graphs.
 Moreover, in cases where LTFT selects the smallest degree nodes less frequently, its performance is worse than Deg-Greedy.
@@
 5.2 Serialization: allows comparing to Deg-Greedy
-Figure 2: The percentage to choose the smallest possible degree node on different part of the (degree-based) serialization.
-We find that the best algorithms (OnlineMIS, ReduMIS, iSCO) and the best performing learning-based algorithm LwD share a similar characteristic pattern, that they have high consistency with degree-based greedy on second and third part of the serialization, while on the first part there is a low consistency.
-On the other hand, PCQO and DIFUSCO has low consistency with degree-based greedy in general.
 In the previous part, we demonstrate that LTFT employs a heuristic similar to Deg-Greedy, selecting the node with the smallest degree in the remaining graph in each round.
@@
 To evaluate the algorithms, we compute the percentage of rounds in which the smallest degree node is selected during serialization, similar to our comparison between LTFT and Deg-Greedy.
-Due to random tie-breaking in Algorithm 1, we repeat the process 100 times and select the serialization with the highest number of rounds selecting the smallest degree node.
-Although Deg-Greedy theoretically achieves 100% smallest-degree selections in its best serialization, random tie-breaking prevents us from perfectly recovering this with 100 repetitions.
-Instead of reporting the overall percentage, we divide the serialization into three equal parts and report the percentages for each part.
-The results, shown in Figure 2, compare the percentage of smallest-degree node selections across different algorithms.
-Due to space constraints, we present results only for ER and BA graphs under selected parameters; additional results are available in Section C.2.
+Due to random tie-breaking in Algorithm 1, we repeat the process 1000 times and select the serialization with the highest number of rounds selecting the smallest degree node.
+In order to provide more fine-grained analysis, instead of reporting the overall percentage, we divide the serialization into three equal parts and report the percentages for each part.
+Although Deg-Greedy theoretically achieves $100\%$ smallest-degree selections in its best serialization, there are exponential number of possible sequences and we cannot perfectly recover this with 1000 repetitions with random tie-breaking for all graphs.
+Nevertheless, for most graphs, Deg-Greedy has $100\%$ or very close to $100\%$ smallest-degree selections in our serialization.
+The results, shown in Figure 1(b), compare the percentage of smallest-degree node selections across different algorithms.
+Due to space constraints, we present results only for ER graphs under selected parameters; additional results for ER graphs and BA graphs are available in Section C.4.
 Algorithms that often perform the best, namely OnlineMIS, ReduMIS, and iSCO, exhibit a consistent pattern after serialization.
@@
 This suggests that while degree-based greedy heuristics may appear shortsighted initially, they are highly effective in the later stages of solution construction.
-Interestingly, LwD, which performs the best among learning-based methods tested in our setting, also shares this pattern.
-As for PCQO and DIFUSCO, they show consistently low percentages of smallest-degree node selections throughout the serialization, particularly on ER graphs.
-Through serialization, we observe distinct differences in node selection patterns among algorithms.
-Our findings suggest that AI-based methods might fail to utilize (PCQO and DIFUSCO) or emphasize too much (LTFT) on simple yet highly effective heuristics, such as greedily selecting the smallest-degree node, which may partly explain their performance limitations.
+Interestingly, LwD and PCQO, which performs worse but close, also shares this pattern, but has higher percentage for the first part, but lower for second and third parts.
+The similar pattern observed among several good-performing algorithms with very different techniques suggest intrinsic similarities among good solutions, regardless the algorithm leads to them.
+Analysis techniques targeting solutions, like serialization here, are especially suitable for discovering characteristics of good solutions, which can provide often overlooked insights for future algorithmic design.
+By contrast, LTFT has very high percentage of smallest-degree nodes selection for all three parts.
+DIFUSCO has low percentage for all three parts.
+Both perform significantly worse than the best-performing algorithms.
+This suggests that these methods might fail to utilize (DIFUSCO) or emphasize too much (LTFT) on simple yet highly effective heuristics, such as greedily selecting the smallest-degree node, which may partly explain their performance limitations.
 In Section C.5, we perform a pseudo-natural serialization for LwD.
-LwD is also a non-backtracking algorithm like LTFT, but it does not have a “natural serialization” like LTFT (Section 4.1) because it chooses several nodes in a step.
+LwD is also a non-backtracking algorithm like LTFT, but it does not have a “natural serialization" like LTFT (Section 5.1) because it chooses several nodes in a step.
 The pseudo-natural serialization performs serialization in each step of LwD.
-The results in Figure 6 align with our “counterfactual” serialization results here.
+The results in Figure 5 align with our “counterfactual" serialization results here.
 5.3 Incorporating local search to improve solution
 In the previous sections, we show that solutions generated by AI-based algorithms generally differ from those produced by degree-based greedy methods, which may explain their inferior performance on MIS problems.
-A natural idea is to enhance these solutions with simple heuristics, such as local search.
-Local search post-processing has also been used for AI-algorithms in previous works (Ahn et al., 2020; Böther et al., 2022).
-We applied the $2$-improvement local search (Andrade et al., 2012) (details in B.7), which is used in KaMIS, as a post-processing step to all algorithms (except OnlineMIS and ReduMIS since they already has local search), and the resulting performance improvements are presented in Table 4.
-Table 4: Incoperating local search as a post-processing procedure We add $2$-improvement local search for the algorithm tested.
+A natural idea to improve them would be to allow post-porcessing their solutions with local search.
+Some previous studies found this beneficial (ahn2020learning, ; boether_dltreesearch_2022, ).
+Table 2 shows the result of applying $2$-improvement local search (andrade2012fast, ), which is used in KaMIS, as a post-processing step for all algorithms (except OnlineMIS and ReduMIS since they already has local search).
+Full results for both ER and BA graphs are in Table 9 and details are in Section B.7.
+Table 2: Incoperating local search as a post-processing procedure on selected ER graphs.
+We add $2$-improvement local search for the algorithm tested.
 We report the performance after the local search and the improvement from local search in ().
-AI-inspired methods still don’t outperform ReduMIS after local search.
-We also observe that if the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search.
-Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search.
-Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 148.50 (5.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 59.00 (3.38) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 453.88 (11.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) BA 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 235.50 (4.50) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 110.25 (2.37) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 723.50 (12.88) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37)
-As shown in Table 4, algorithms like PCQO and DIFUSCO benefit significantly more from the local search post-processing compared to others, such as Deg-Greedy and LwD.
+Full results for ER and BA graphs in Table 9.
+Heuristics GPU-acc Learning-based $(n,d)$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 159.13 (0.5) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 60.88 (0.75) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 469.75 (5.5) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37)
+As shown in Table 2, DIFUSCO benefit significantly more from the local search post-processing compared to others, such as Deg-Greedy and LwD.
 This observation aligns with our earlier findings: If the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search.
 Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search.
-In addition, although all algorithms except iSCO have improvements after local search, they still perform worse than ReduMIS in most cases.
-In summary, our analysis highlights a promising direction for designing machine learning-based combinatorial optimization algorithms.
-Rather than relying solely on end-to-end methods like PCQO or DIFUSCO, incorporating classical heuristics, such as greedily selecting the smallest-degree node, into the overall algorithm may yield better results.
+In addition, although all algorithms except iSCO have improvements after local search, they still perform worse than ReduMIS.
+6 Call for New Directions: Issues and Solutions
+Building on our empirical findings and analysis on MIS as a testbed, we highlight several critical issues in the current research landscape of AI-inspired methods for combinatorial optimization (CO), and propose actionable directions to address them.
+We expect these issues and solutions are relevant for a wide range of CO problems, given similar challenges observed in tasks like TSP (joshi2021learning, ; sun2023difusco, ; xia2024position, ) and theoretical barriers identified for GNN-based approaches (gamarnik2023barriers, ).
+The following issue-solution pairs aim to guide future work toward more robust, insightful, and effective progress in this field.
+Issue 1: Limited benchmarks and biased evaluation
+Recent research on AI-inspired methods for combinatorial optimization (CO) often relies on limited test datasets and focuses on small-scale instances with tight inference time constraints, which can artificially favor AI-based methods (boether_dltreesearch_2022, ; ahn2020learning, ; sun2023difusco, ).
+While state-of-the-art classical solvers typically require more time, simple heuristics (like degree-based greedy for MIS), which are fast and possibly competitive on small instances, are often overlooked.
+Furthermore, current studies overlook that AI-inspired methods, especially on larger graphs, demand much higher computational resources—often exceeding 80GB of GPU memory for ER graphs with $10000$ nodes (Table 1), making them infeasible without multi-GPU setups.
+By contrast, KaMIS runs on a single CPU thread, scales efficiently up to millions of nodes (Table 8), and can be parallelized.
+Thus, classical solvers are often more scalable and practical than suggested by current benchmarks, which may not reflect true algorithmic strengths and weaknesses.
+Previous works (joshi2021learning, ; sun2023difusco, ; xia2024position, ) have highlighted problematic evaluations of neural-based methods on TSP, suggesting similar issues throughout CO.
+Solution: Rigorous and comprehensive benchmarking
+A critical step forward is to benchmark AI-inspired methods against classical solvers on a wider range of CO problems across varied instance sizes and realistic runtime settings.
+Comprehensive benchmarks, like (boether_dltreesearch_2022, ), as well as detailed experiments (Section 4) and analyses (Section 5) in our paper, are necessary for meaningful comparisons.
+Future evaluations should also report computation resources used (e.g., GPU memory) and allow for flexible inference times, ensuring a fair assessment across methods.
+Issue 2: Intrinsic hardness and limits of learning
+Our experiments (Section 4) show that learning-based methods consistently underperform leading classical solvers, even on in-distribution data thought to be ideal for ML.
+This suggests that problems like MIS may be fundamentally hard for current learning paradigms.
+Theoretical works highlight barriers for local and GNN-based algorithms, such as the Overlap Gap Property (OGP) (gamarnik2014limits, ; gamarnik2021overlap, ; gamarnik2023barriers, ).
+Solution: Theoretical and empirical understanding of limitations
+Addressing these limitations requires a combination of theoretical and empirical approaches.
+Future research should deepen understanding of theoretical barriers like OGP and develop empirical tools, such as serialization analysis (Section 5), to pinpoint failure modes of AI-inspired methods.
+This dual approach will clarify what is fundamentally difficult for learning in CO and guide new algorithmic developments.
+Issue 3: Insufficient use and understanding of classical heuristics
+Current AI-inspired algorithms for CO problems either attempt to fully replace classical heuristics or substitute classical components with learned models, often without understanding why classical methods succeed.
+Previous works (boether_dltreesearch_2022, ; xia2024position, ) reveal that the seeming success of some hybrid approaches, including neural network-guided tree search for MIS (li2018combinatorial, ) and TSP (xia2024position, ), typically derive most of their gains from embedded classical reductions rather than from the learned components.
+Solution: Learning from and integrating classical heuristics
+Progress demands that researchers move beyond treating classical solvers as black-box baselines.
+Instead, AI researchers tackling CO problems should analyze and learn the effective components of classical heuristics—such as reduction (xiao2013confining, ; akiba2016branch, ), local search (andrade2012fast, ), and evolutionary algorithms (lamm2015graph, ).
+It is also important to use empirical ablation studies to identify key ingredients, as done in (boether_dltreesearch_2022, ) and Section 5.3.
+Results in Section 5.2 show similarities among good solutions obtained from vastly differnt algorithms.
+Furthermore, our analysis Section 5 highlights that rather than relying solely on end-to-end methods like DIFUSCO, integrating classical heuristics such as greedily selecting the smallest-degree node into the overall algorithm may yield better results.
 One potential approach could involve using machine learning algorithms to identify a small subset of nodes, followed by a degree-based greedy method to complete the solution.
-4.4 KaMIS refutes famous conjecture?
-For theoretical analysis for MIS on ER graphs and regular graphs, see Coja-Oghlan & Efthymiou (2015); Wormald et al.
-(1999); Barbier et al.
-(2013); Gamarnik & Sudan (2014).
-In ER graphs with $n$ nodes and average degree $d$, the MIS has size $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$, and simplest random greedy achieves half-optimal at $\frac{n\ln d}{d}$ (Grimmett & McDiarmid, 1975).
-Yet, there is no known polynomial-time algorithm which can achieve MIS size $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ and it is conjectured that polynomial-time algorithms cannot do better than $(1+o(1))\frac{n\ln d}{d}$ (Coja-Oghlan & Efthymiou, 2015).
-Thus it is natural to measure the goodness of an algorithm by the ratio of the MIS size obtained to $\frac{n\ln d}{d}$.
-This ratio can faciliate comparison across different $n$ and $d$’s.
-Figure 3 plots this ratio for several algorithms.
-Surprisingly for ReduMIS and OnlineMIS this ratio is consistently larger than $1.2$ and often larger than $1.3$ even for fairly large $n,d$.
-These results seem to bring the conjecture in doubt, but not disprove per se since the conjecture concerns asymptotically large $n$ and $d$.
-But our results could encourage further analysis and potential collaboration between researchers on theoretical and empirical aspects of MIS on random graphs.
-Figure 3: Heatmap for ratios of MIS size to $\frac{n\ln d}{d}$ on ER graphs.
-We find that the ratio for ReduMIS and KaMIS is consistently larger than 1.2, suggesting that ReduMIS and KaMIS might surpass the conjectured upper bound (Coja-Oghlan & Efthymiou, 2015).
-5 Related Works
-Classical and heuristic methods for MIS
-Classical methods for MIS range from simple greedy algorithms to advanced solvers like KaMIS which involves a number of heuristics.
-There are various existing heuristics for MIS, such as reduction techniques (Butenko et al., 2002; Xiao & Nagamochi, 2013; Akiba & Iwata, 2016), local search (Andrade et al., 2012), and evolutionary algorithms (Back & Khuri, 1994; Borisovsky & Zavolovskaya, 2003; Lamm et al., 2015).
-KaMIS (Lamm et al., 2017; Dahlum et al., 2016; Hespe et al., 2019) was developed based on these heuristics.
-In addition, MIS can be formulated into a binary integer programming problem (Nemhauser & Trotter Jr, 1975), which can be solved by the state-of-art integer programming solver Gurobi Gurobi Optimization, LLC (2024).
-Machine learning for combinatorial optimization
-In recent years, various ML-based algorithms have been developed for the MIS problem and most of them are based on graph neural networks (GNNs).
-Some of them using supervised learning (Li et al., 2018; Sun & Yang, 2023; Li et al., 2024) and requires labelling training data using classical solvers.
-Alternatively, those based on reinforcement learning (Khalil et al., 2017; Ahn et al., 2020; Qiu et al., 2022; Sanokowski et al., 2023) and other unsupervised learning objectives (Karalias & Loukas, 2020; Sun et al., 2022; Zhang et al., 2023; Sanokowski et al., 2024) do not require labeled training data.
-Zhang et al.
-(2023) (LTFT) uses GFlowNets (Bengio et al., 2021a) which is related to reinforcement learning.
-Notably, Ahn et al.
-(2020); Sanokowski et al.
-(2023); Zhang et al.
-(2023) model MIS problem as a Markov Decision Process (MDP) and generate the solution step-by-step (autoregressively).
-While Sanokowski et al.
-(2023) fixes the order of node updates, Ahn et al.
-(2020) (LwD) and Zhang et al.
-(2023) (LTFT) choose which node to update at each step so that they have a “natural” or “pseudo-natural” serialization and most suitable for our analysis in Section 4.1 and Section C.3.
-In addition, Sun & Yang (2023) (DIFUSCO) and Sanokowski et al.
-(2024) (DiffUCO) both utilizes diffusion model.
-Most of the algorithms above also work on other types of graph CO problems, such as Maximum Cut, Maximum Vetex Cover, and Minimum Dominating Set.
-Some (Khalil et al., 2017; Qiu et al., 2022; Sun & Yang, 2023) also work for the Travelling Salesman Problem (TSP).
-Other GPU-based solvers for CO
-Recently, some non-learning GPU-based solvers for CO problems have been developed.
-Sun et al.
-(2023) developed a GPU-accelerated sampling based method which works on MIS, Max Cut, and TSP.
-Schuetz et al.
-(2022); Ichikawa (2023) uses GNNs conduct non-convex optimization for MIS without machine learning.
-Alkhouri et al.
-(2024) uses direct quadratic optimization without GNNs for MIS.
-Benchmarks for MIS
-Böther et al.
-(2022) provides a benchmark for several MIS algorithms including Gurobi (Gurobi Optimization, LLC, 2024), KaMIS (ReduMIS(Lamm et al., 2017), Intel-Treesearch (Li et al., 2018), DGL-Treesearch, and LwD Ahn et al.
-(2020).
-It is the only MIS benchmark we know about including recent AI-inspired method, though it only focuses its comparsion for learning-based tree search algorithms.
-It suggests that LwD is better than learning-based tree search algorithms.
-This aligns with our results where LwD performs the best among learning-based algorithms.
-This benchmark covers various types of random graphs and several real-world datasets, so it is a good reference benchmark for comparison over different types of graphs.
-Though unlike our benchmark, it does not provide comparison across various size and density for random graphs.
-Our benchmark fills this gap and provides a more detailed comparison.
-We also include many newer AI-inspired algorithms, and greedy algorithms which leads to detailed analysis like serialization.
-Angelini & Ricci-Tersenghi (2023) provides a comparison between Schuetz et al.
-(2022) and Deg-Greedy over regular graphs.
-The algorithm papers also report experiments comparing with previous algorithms, but those comparisons usually only focus on a few datasets and a few selected baselines.
-6 Conclusion and Takeaways
-Given the great interest in designing “general purpose AI reasoners”, it is interesting to check how well recent AI-based methods have fared in combinatorial optimization, a field with a long history of ingenious hand-designed algorithms.
-Our careful empirical comparisons of such AI-inspired methods with classical methods on MIS problem showed that none of the new methods outperform ReduMIS, the best CPU-based MIS solver, which builds up the independent set iteratively, sometimes backtracking (i.e.
-delete a vertex from the current set).
-As the graphs get larger or denser, the superiority of ReduMIS becomes more evident, whereas several AI-inspired algorithms drop to performing no better than trivial classical algorithms such as Deg-Greedy.
-Further analysis shows that the fact that AI-inspired algorithms like DIFUSCO and PCQO select the independent set in one shot may be handicapping themselves by foregoing the benefits of local search.
-Methods like Deg-Greedy and LTFT select one node at a step and do not backtrack end up performing weakly.
-Interestingly, the best-performing learning-based algorithm LwD is doing something in the middle: selecting several nodes at a time, which may allow a more effective approximation to local search.
-Using local search as post-processing improves the solutions of AI-inspired methods more than the CPU-based algorithms such as degree-based greedy.
-Despite that, with the local search step, these AI-inspired methods still perform worse than ReduMIS.
-This suggests that a more promising method to use AI and machine learning techniques for combinatorial optimization might be to teach the model the value of local search instead of simply use it as a post-processing.
-Acknowledgement
-YW, HZ, and SA acknowledge funding from DARPA. The authors would like to thank Alvaro Velasquez, Ismail Alkhouri, Kaifeng Lyu, Sadhika Malladi, and Pravesh Kothari for helpful discussions.
-Appendix
+These insights should then be systematically incorporated into AI-inspired frameworks, enabling principled hybrid algorithms that embed classical heuristics as core components.
+7 Alternative Views
+While our paper primarily targets algorithm designers and emphasizes the need to rethink and design better AI-inspired methods for combinatorial optimization problems, users who need to decide which method to use may hold different perspectives.
+Although our analysis clearly shows that AI-inspired methods currently lag behind classical heuristic solvers like KaMIS in terms of solution quality, there are situations where AI-inspired methods may be preferable.
+Learning-based AI methods often offer shorter inference times and can achieve comparable, or occasionally better, results when computation time is limited (boether_dltreesearch_2022, ; ahn2020learning, ; zhang2023let, ; alkhouri2024dataless, ).
+In scenarios where speed is prioritized over optimality, or where strict inference-time constraints exist, current AI-inspired approaches may be preferable options.
+Nevertheless, it is still important to compare these methods against classical fast heuristics in time-sensitive settings.
+For example, in the MIS problem, the simple degree-based greedy algorithm also provides fast computation and often competitive performance, as shown in Section 4, and should be considered a strong baseline.
+Additionally, classical methods like KaMIS and degree-based greedy typically require fewer computational resources, such as GPU memory, than AI-inspired methods.
+Ultimately, there is no single “best” algorithm for the MIS problem—or for other combinatorial optimization problems—when practical use cases are considered.
+Users should select methods based on their specific requirements and constraints.
+Appendices
 Appendix A More Related Works
 More on theoretical results for MIS
-For random ER graphs with number of nodes $n$ and average degree $d$, the upper bound of MIS size is $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$(Coja-Oghlan & Efthymiou, 2015).
-Grimmett & McDiarmid (1975) proves that the simplest Ran-Greedy can achieve half-optimal at $\frac{n\ln d}{d}$ on random ER graphs.
-Despite that, there is no known existing polynomial algorithm which can reach MIS size of $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ for asymptotically large $n$ and $d$(Coja-Oghlan & Efthymiou, 2015).
-Coja-Oghlan & Efthymiou (2015) also suggests that the reason is likely independent sets of size larger than $(1+{\epsilon})\frac{n\ln d}{d}$ forms an intricately ragged landscape, where local algorithms will stuck.
-Gamarnik & Sudan (2014); Rahman & Virág (2017) proves that for local algorithms (which is defined to only use information from a constant neighborhood of a node to decide whether the node is in the independent set) are at most half optimal for independent set on random $d$-regular graphs.
-Gamarnik & Sudan (2014) suggests this is due to a property of the MIS problem, which they denote as the Overlap Gap Property (OGP) (Gamarnik, 2021).
-Gamarnik (2023) suggests that graph neural networks (GNNs) are also a type of local algorithms and thus being limited by OGP.
-While most AI-inspired MIS algorithms use GNN, the proof only applies to algorithms use GNNs as the only component to find solutions like (Schuetz et al., 2022), so it may not apply directly to more complicated algorithms like those tested in our paper.
+For random ER graphs with number of nodes $n$ and average degree $d$, the upper bound of MIS size is $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$[51].
+[52] proves that the simplest Ran-Greedy can achieve half-optimal at $\frac{n\ln d}{d}$ on random ER graphs.
+Despite that, there is no known existing polynomial algorithm which can reach MIS size of $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ for asymptotically large $n$ and $d$[51].
+[51] also suggests that the reason is likely independent sets of size larger than $(1+{\epsilon})\frac{n\ln d}{d}$ forms an intricately ragged landscape, where local algorithms will stuck.
+[49, 53] proves that for local algorithms (which is defined to only use information from a constant neighborhood of a node to decide whether the node is in the independent set) are at most half optimal for independent set on random $d$-regular graphs.
+[49] suggests this is due to a property of the MIS problem, which they denote as the Overlap Gap Property (OGP) [50].
+[25] suggests that graph neural networks (GNNs) are also a type of local algorithms and thus being limited by OGP.
+While most AI-inspired MIS algorithms use GNN, the proof only applies to algorithms use GNNs as the only component to find solutions like [23], so it may not apply directly to more complicated algorithms like those tested in our paper.
 Yet, it may still suggests a reason why GNN-based algorithms (including most AI-inspired algorithms) cannot outperform classical heuristics like KaMIS.
-In addition, Barbier et al.
-(2013) provides a conjectured tighter upper bound than $\frac{2n\ln d}{d}$ for $d$-regular graphs using the hard-core model in physics.
-Ding et al.
-(2016) proves a similar tighter upper bound for $d$-regular graphs.
-Wormald (2003) gives average-case performance for Deg-Greedy on $d$-regular graphs.
+In addition, [54] provides a conjectured tighter upper bound than $\frac{2n\ln d}{d}$ for $d$-regular graphs using the hard-core model in physics.
+[55] proves a similar tighter upper bound for $d$-regular graphs.
+[56] gives average-case performance for Deg-Greedy on $d$-regular graphs.
 More on classical heuristics
 Over the past few decades, significant progress has been made in tackling NP-hard combinatorial optimization (CO) problems by developing approximation algorithms and heuristic methods.
-Approximation algorithms provide provable guarantees on solution quality and have led to groundbreaking results for classical problems, such as the Maximum Independent Set (MIS), Traveling Salesperson Problem (TSP), and Maximum Cut (Boppana & Halldórsson, 1992; Laporte, 1992; Goemans & Williamson, 1995).
+Approximation algorithms provide provable guarantees on solution quality and have led to groundbreaking results for classical problems, such as the Maximum Independent Set (MIS), Traveling Salesperson Problem (TSP), and Maximum Cut [57, 58, 59].
 As we mentioned in Section 2, there are various existing heuristics for MIS.
 For exmaple, reduction techniques reduce the graph into smaller instances.
-Akiba & Iwata (2016) and Xiao & Nagamochi (2013) have shown a variety of reduction techniques which work well for MIS problem.
-Butenko et al.
-(2002); Bourgeois et al.
-(2012) uses reduction techniques to develop efficient exact algorithms for MIS.
+[32] and [31] have shown a variety of reduction techniques which work well for MIS problem.
+[30, 60] uses reduction techniques to develop efficient exact algorithms for MIS.
 Local search improves an existing independent set by removing a small number of nodes and insert other eligible nodes.
-Andrade et al.
-(2012) gives an efficient local search algorithm for MIS and has been used as a subprocess for several MIS solvers.
+[33] gives an efficient local search algorithm for MIS and has been used as a subprocess for several MIS solvers.
 Evolutionary algorithms combine several existing solutions into a new solution.
-Examples include Back & Khuri (1994); Borisovsky & Zavolovskaya (2003); Lamm et al.
-(2015).
-KaMIS (Lamm et al., 2017; Dahlum et al., 2016) was developed based on many of these techniques above.
-In addition, the MIS problem can also be relaxed into semi-definite programming (SDP), which leads to several approximation algorithms (Halperin, 2002; Bansal, 2014).
+Examples include [34, 35, 36].
+KaMIS [28, 37] was developed based on many of these techniques above.
+In addition, the MIS problem can also be relaxed into semi-definite programming (SDP), which leads to several approximation algorithms [61, 62].
 More on AI methods for combinatorial optimization
-In addition to Section 5, we note that Sun et al.
-(2022); Sanokowski et al.
-(2023; 2024) use annealing techniques.
-Sun & Yang (2023) was developed based on Qiu et al.
-(2022), and Sanokowski et al.
-(2024) was based on Sanokowski et al.
-(2023).
-Sanokowski et al.
-(2024) can also be considered as an extension of Karalias & Loukas (2020).
-Besides MIS, Maximum Cut, and TSP, people also considered to use AI-Inspired methods for other combinatorial optimization problems, including Vehicle Routing Problems (including TSP) (Kool et al., 2018; Chen & Tian, 2019; Delarue et al., 2020; Li et al., 2021; Zheng et al., 2021; Ye et al., 2024), Job Scheduling Problems (Lin et al., 2019; Baer et al., 2019; Zhang et al., 2020; Ye et al., 2024), Boolean Satisfiability (SAT) (Amizadeh et al., 2019; You et al., 2019; Kurin et al., 2020; Li & Si, 2022; Li et al., 2023), and Casual Discovery (Zheng et al., 2018; Zhu et al., 2020; Sanchez et al., 2022).
+In addition to Section 2, we note that [43, 41, 16] use annealing techniques. [15] was developed based on [40], and [16] was based on [41]. [16] can also be considered as an extension of [42].
+Besides MIS, Maximum Cut, and TSP, people also considered to use AI-Inspired methods for other combinatorial optimization problems, including Vehicle Routing Problems (including TSP) [8, 63, 64, 65, 66, 67], Job Scheduling Problems [68, 69, 70, 67], Boolean Satisfiability (SAT) [71, 72, 73, 74, 75], and Casual Discovery [76, 77, 78].
 Appendix B Detailed Experiment Setup
@@
 Erdős-Renyí (ER) graph
-ER graphs (Erdös & Rényi, 1959) have 2 parameters: the number of nodes $n$ and the average degree $d$.
+ER graphs [45] are random graphs where edges are connected uniformly at random (with a fixed probability or number of edges).
+We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing number of edges at $\frac{nd}{2}$.
+ER graphs are fundamental objects in network sciences [79] and random graph theory [80].
+There are also theoretical analysis and conjecture upper bound for MIS on ER graphs [51].
+Previous works [24, 7, 15, 21, 18] used it as test graphs for MIS, though without varying parameters as we did.
+ER graphs have 2 parameters: the number of nodes $n$ and the average degree $d$.
 For graphs with given $(n,d)$ We generate them by choosing $M=\frac{nd}{2}$ edges uniformly at random between $n$ nodes.
@@
 Barabási–Albert (BA) graph
-BA graphs (Albert & Barabási, 2002) also have 2 parameters: the number of nodes $n$ and the generation parameter $m$, with $n>m$.
+Barabási–Albert (BA) graphs [46] are random graphs generated by a probabilistic growth process, whereby new nodes preferentially attach to existing nodes with higher degrees, mimicking real-world networks such as Internet, citation networks, and social networks [46, 81].
+For BA graphs, we vary $2$ parameters: number of nodes $n$ and parameter $m$ (not number of edges).
+The average degree of BA graphs can be approximated as $2m$.
+BA graphs also have 2 parameters: the number of nodes $n$ and the generation parameter $m$, with $n>m$.
 For a given $(n,m)$, the generation process initializes with $m$ nodes, and then add 1 node at each step.
@@
 For the main experiments, similar to ER graphs, we also use $n=\{100,300,1000,3000,10000\}$. Since the average degree $d\approx 2m$, we use $m=\{5,15,50,150,500,1500\}$ with $2m<n$.
+RB graphs
+RB graphs are derived from Model RB [47], a random constraint satisfaction problem (CSP) model.
+RB graphs are considered difficult instances for MIS due to their structured randomness and high solution hardness.
+We use two datasets from [21] (also used in [16]), RB-small (200–300 nodes) and RB-large (800–1200 nodes), to benchmark learning-based solvers.
 Real-world graphs
+We pick REDDIT-MULTI-5K and COLLAB [48] from TUDataset website [82], since they have enough graphs for training and graph sizes not too small.
+REDDIT-MULTI-5K has $508.52$ average nodes and $594.87$ average edges.
+They are mostly very sparse graphs.
+COLLAB has $74.49$ average nodes and 2457.78 average edges.
+They are mostly small but dense graphs.
 Since we need at least $4100$ graphs to train and test our algorithms and the graphs should not be too small, it is difficult to find such datasets.
-Fortunately, Morris et al.
-(2020) provides a website www.graphlearning.io which includes many graph datasets prepared by them or collected from other works.
-Although most of the datasets are still too small or having too small graphs, we are able to find 2 datasets: REDDIT-MULTI-5K and COLLA, both from Yanardag & Vishwanathan (2015).
+Fortunately, [82] provides a website www.graphlearning.io which includes many graph datasets prepared by them or collected from other works.
+Although most of the datasets are still too small or having too small graphs, we are able to find 2 datasets: REDDIT-MULTI-5K and COLLA, both from [48].
 REDDIT-MULTI-5K has $4999$ graphs with average nodes $508.52$ and average edges $594.87$, so the graphs are generally very sparse.
@@
 We were not able to find real-world datasets with enough size of more larger and denser graphs, which are generally more difficult for MIS algorithms.
-The dataset DIMACS used in Böther et al.
-(2022) contains such graphs but they only have $37$ graphs which is not enough for training.
+The dataset DIMACS used in [24] contains such graphs but they only have $37$ graphs which is not enough for training.
 B.3 Hardware configurations
@@
 We did not include results of Ran-Greedy in the main paper because its performance is significantly lower than all other algorithms.
-It can be considered as a baseline and has theoretical significance, since it has provable guarantee for random graphs (Grimmett & McDiarmid, 1975).
+It can be considered as a baseline and has theoretical significance, since it has provable guarantee for random graphs [52].
 In order to match the best-of-20 sampling we used for the learning-based-algorithms, we also ran Deg-Greedy for $20$ times and report the best results in the main experiments ($n\leq 10000$).
@@
 KaMIS (OnlineMIS and ReduMIS)
-KaMIS (Karlsruhe Maximum Independent Sets) (Lamm et al., 2017) (https://karlsruhemis.github.io/) is the state-of-art heuristic solver for MIS and has been used as a baseline in many previous works.
-It provides 2 algorithms for the MIS problem: ReduMIS (Lamm et al., 2017) and OnlineMIS (Dahlum et al., 2016).
+KaMIS (Karlsruhe Maximum Independent Sets) [28] (https://karlsruhemis.github.io/) is the state-of-art heuristic solver for MIS and has been used as a baseline in many previous works.
+It provides 2 algorithms for the MIS problem: ReduMIS [28] and OnlineMIS [37].
 We can provide a time-limit for both algorithm.
@@
 iSCO
-iSCO (improved Sampling for Combinatorial Optimization) (Sun et al., 2023) is a GPU-accelerated sampling-based method.
+iSCO (improved Sampling for Combinatorial Optimization) [17] is a GPU-accelerated sampling-based method.
 It does not require learning.
-According to (Sun et al., 2023), the main benefit is that it can process a large batch of graphs in parallel thus improve efficiency.
+According to [17], the main benefit is that it can process a large batch of graphs in parallel thus improve efficiency.
 While processing a small number of graphs like in our case (8 test graphs), it still requires significant time, often longer than ReduMIS.
-We use the code from the codebase of DISCS (Goshvadi et al., 2024), which is a follow-up paper for iSCO, as iSCO did not provide the codebase.
+We use the code from the codebase of DISCS [83], which is a follow-up paper for iSCO, as iSCO did not provide the codebase.
 We use 1 80GB $A100$ GPU to run iSCO with all test graphs together.
@@
 PCQO
-PCQO (Parallelized Clique-Informed Quadratic Optimization) (Alkhouri et al., 2024) is a GPU-accelerated gradient-based optimization algorithm, which directly optimize on the quadratic loss function of the MIS problem.
+PCQO (Parallelized Clique-Informed Quadratic Optimization) [18] is a GPU-accelerated gradient-based optimization algorithm, which directly optimize on the quadratic loss function of the MIS problem.
 The loss function for a graph $G=(V,E)$ and solution vector ${\bm{x}}$ is
@@
 for each $(n,d)$ pair in ER graphs and each $(n,m)$ pair in BA graphs).
-Although Alkhouri et al.
-(2024) reported that using the third term in the loss function (i.e.
-nonzero $\gamma^{\prime}$) can lead to better answers comparing to only using 2-term loss function (set $\gamma^{\prime}=0$), we found that a nonzero $\gamma^{\prime}$ makes the algorithm more sensitive to hyperparameters and difficult to perform hyperparameter tuning for our benchmark with many datasets.
-Therefore, we only use $2$-term loss function by setting $\gamma^{\prime}=0$.
-In the grid search for hyperparameters, we use learning rate $0.001$ (default), momentum $\{0.5,0.9\}$, $\gamma\in\{4,20,100,500,2500\}$, and report the results obtained from the best hyperparameters.
+We use the hyperparamter search range suggested by the authors [18].
+We use learning rate $\{0.0001,0.00001,0.000001\}$, momentum $0.9$, $\gamma\in\{200,500,1000,2000,5000\}$, $\gamma^{\prime}=1$, batch size $256$, number of steps $27000$, steps per batch $450$.
+We then report the results obtained from the best set of hyperparameter.
 However, for $n\geq 30000$, the grid search within this domain does not provide solution with reasonable size (worse than Ran-Greedy), so we do not report results for larger graphs.
 Despite that, there may exist a better set of hyperparamters which could make this algorithm perform well on larger graphs.
+Automatic hyperparameter search could significantly improve the usability of this algorithm.
 We use 1 80GB $A100$ GPU for all test graphs together with a time limit of 96hrs. It is able to run all experiments from $n\leq 10000$, and often requires shorter time compared to iSCO and KaMIS.
@@
 LwD
-LwD (Learning what to Defer) (Ahn et al., 2020) is a reinforcement learning based algorithm which requires training data to learn the policy.
+LwD (Learning what to Defer) [7] is a reinforcement learning based algorithm which requires training data to learn the policy.
 It models the MIS problem as a Markov Decision Process (MDP), where in each step it selects some (possibly 00, $1$, or multiple) nodes to add into the independent set.
 It is a non-backtracking algorithm as the added nodes are never taken out.
-Böther et al.
-(2022) also included this algorithm in their benchmark.
-We use the code from Böther et al.
-(2022) since it provides better functionality than the original codebase.
-We use the default setting provided by Böther et al.
-(2022) in their MIS Benchmark codebase (https://github.com/MaxiBoether/mis-benchmark-framework), but change the number of samples to $20$ (default is $10$) for test sampling, in order to match the best-of-20 sampling in our benchmark.
+[24] also included this algorithm in their benchmark.
+We use the code from [24] since it provides better functionality than the original codebase.
+We use the default setting provided by [24] in their MIS Benchmark codebase (https://github.com/MaxiBoether/mis-benchmark-framework), but change the number of samples to $20$ (default is $10$) for test sampling, in order to match the best-of-20 sampling in our benchmark.
 We use 1 80GB $A100$ GPU to train for each datasets and test with the same GPU.
@@
 LTFT
-LTFT (Let the Flows Tell) (Zhang et al., 2023) similar to LwD, also model the MIS problem as a MDP and non-backtracking.
+LTFT (Let the Flows Tell) [21] similar to LwD, also model the MIS problem as a MDP and non-backtracking.
 The difference is that it only choose $1$ node at each step, making it more similar to Deg-Greedy.
-The node chosen at each step is chosen by a GFlowNet (Bengio et al., 2021a), which is trained by in-distribution training data.
-We use the default setting provide by Zhang et al. (2023) for training with $20$ epochs. By default setting, it has best-of-20 sampling and report the best solution found.
+The node chosen at each step is chosen by a GFlowNet [19], which is trained by in-distribution training data.
+We use the default setting provide by [21] for training with $20$ epochs. By default setting, it has best-of-20 sampling and report the best solution found.
 We use 1 80GB $A100$ GPU to train for each datasets and test with the same GPU.
@@
 DIFUSCO
-DIFUSCO (Diffusion Solvers for Combinatorial Optimization) (Sun & Yang, 2023) trains a diffusion model using supervised learning to produce a solution for the MIS.
+DIFUSCO (Diffusion Solvers for Combinatorial Optimization) [15] trains a diffusion model using supervised learning to produce a solution for the MIS.
 The diffusion model provides an entire solution so it is a one-shot algorithm.
@@
 For graphs we used for training ($n\leq 3000$), ReduMIS gives the same performance compared to a time limit of 24hrs.
-We use the default setting in (Sun & Yang, 2023) except that we use $50$ diffusion steps throughout training and testing, and $20$ samples for testing to be aligned with best-of-20 sampling in other methods.
+We use the default setting in [15] except that we use $50$ diffusion steps throughout training and testing, and $20$ samples for testing to be aligned with best-of-20 sampling in other methods.
 We train the model for $50$ epochs (default) for each dataset.
@@
 We report results where the training can be completed.
-Although Sun & Yang (2023) suggested that DIFUSCO has some generalization ability.
+Although [15] suggested that DIFUSCO has some generalization ability.
 We found the performance degrade significantly for out-of-distribution trained models (specifically trained on smaller graphs with the same average degree but test on larger graphs), we did not report the results of larger graphs where the in-distribution training cannot finish.
 DiffUCO
-DiffUCO (Diffusion for Unsupervised Combinatorial Optimization) (Sanokowski et al., 2024) is also a diffusion model based algorithm but unlike DIFUSCO it uses unsupervised learning.
+DiffUCO (Diffusion for Unsupervised Combinatorial Optimization) [16] is also a diffusion model based algorithm but unlike DIFUSCO it uses unsupervised learning.
 The diffusion model is trained to sample the solution of low energy state.
 It also provides an entire solution and is also a one-shot algorithm.
-We use the default setting in Sanokowski et al.
-(2024) for RB-large MIS task (in their Appendix C.5).
+We use the default setting in [16] for RB-large MIS task (in their Appendix C.5).
 During testing, we use conditional expectation with $20$ samples to align with best-of-20 sampling in other algorithms.
@@
 It can be used as a post-processing technique, or be used as sub-procedures in more complicated algorithms like KaMIS.
-Andrade et al.
-(2012) provides an efficient local search algorithm.
-Part of it is to find $2$-improvement, which is the part used as sub-procedure in KaMIS (Lamm et al., 2017; Dahlum et al., 2016).
+[33] provides an efficient local search algorithm.
+Part of it is to find $2$-improvement, which is the part used as sub-procedure in KaMIS [28, 37].
 The local search algorithm for $2$-improvement for a given independent set $I$ is as follows.
@@
 If no improvement is found, reinsert $x$ into $S$ to turn it back to $I$.
-Every vertex is scanned $O(1)$ times in this algorithm so it can find a $2$-improvement (if there exists) in $O(m)$ time according to Andrade et al.
-(2012).
+Every vertex is scanned $O(1)$ times in this algorithm so it can find a $2$-improvement (if there exists) in $O(m)$ time according to [33].
 We implemented this algorithm in Python and use it as a post-processing for the solutions produced by the algorithm we test.
@@
 In this section, we show more experiment results.
-Section C.1 shows more experiment results on much larger graphs, where the AI-inspired methods cannot handle.
+Section C.1 shows the performance of various algorithms as in Section 4 on other types graphs.Section C.2 shows more experiment results on much larger graphs, where the AI-inspired methods cannot handle.
 Section C.4 show the serialization results on more graphs.
@@
 Section C.6 shows the full results when adding local search as a post-processing procedure.
-Finally, Section C.5 discusses more results on the ratio of the solution found by different algorithms to $\frac{n\ln d}{d}$ on ER graphs, which relates to the theoretical conjecture mentioned in Section 4.4.
-C.1 Larger graphs
+C.1 Results for BA graphs, RB graphs, and real-world graphs
+In this section we report the performance of various algorithms on Barabási–Albert (BA) graphs in Table 5, RB graphs in Table 6, and real-world graphs in Table 7.
+[…]
+Heuristics GPU-acc Learning-based $n$ $m$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 5 39.50 39.50 39.50 39.50 39.50 39.50 38.12 39.38 39.25 15 21.00 21.63 21.63 21.62 21.63 21.62 20.62 21.25 21.25 300 5 121.88 123.13 123.13 123.12 123.00 123.12 115.62 122.88 123.00 15 69.38 71.38 71.38 71.38 71.25 70.75 64.75 69.50 69.25 50 38.00 49.88 50.00 50.00 50.00 50.00 41.75 50.00 50.00 1000 5 412.63 417.13 417.13 417.12 415.75 416.00 385.75 417.12 416.00 15 232.75 245.00 246.38 246.25 242.00 243.12 224.25 236.38 240.75 50 107.50 115.75 116.88 116.75 113.50 113.00* 106.38 105.25 56.38† 150 81.50 150.00 150.00 150.00 150.00 150.00* 82.25 135.00† 3000 5 1236.63 1257.00 1257.13 1255.62 1243.00 1248.12 1177.25 1254.38 1252.63† 15 709.50 749.63 754.50 752.00 722.25 727.75* 661.88 726.38 737.38† 50 335.38 362.63 369.75 368.25 357.13 336.88* 323.88 142.25† 150 147.25 160.25 165.75 164.00 152.75 144.75 500 141.00 500.00 500.00 491.00 223.88 10000 5 4128.75 4206.00 4205.38 4190.47 3640.75 15 2377.88 2525.00 2534.00 2517.07 1946.38 50 1122.13 1228.88 1251.13 1241.98 938.38 150 511.25 555.75 580.13 573.37 433.88 500 192.50 209.13 217.50 150.63 1500 67.25 1500.00 1500.00
+Table 6: (Comparison of the performance of different algorithms on RB graphs) Table 1 on RB graphs [47]. Datasets are from [21]. The numbers within ±$\pm 1\%$ of the best are highlighted.
+Dataset no.
+of nodes Heuristics GPU-acc Learning-based Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO RB-small 200-300 19.11 20.11 20.14 20.05 20.08 16.44 18.63 18.23 19.26 RB-large 800-1200 39.16 42.66 42.95 41.53 39.62 32.55 38.32 36.22 38.91
+Table 7: (Comparison of the performance of different algorithms on real-world graphs) Table 1 on real-world datasets.
+In general, graphs in REDDIT-MULTI-5K are very sparse, while COLLAB are dense but small.
+The numbers within ±$\pm 1\%$ of the best are highlighted.
+Dataset Heuristics GPU-acc Learning-based Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO REDDIT-MULTI-5K 350.73 350.73 350.66 350.73 344.47 350.73 343.35 350.72 350.69 COLLAB 8.68 8.70 8.70 8.70 8.57 8.69 8.70 8.75 8.70
+C.2 Larger graphs
 Table 8 reports our results for large ER graphs not reported in Table 1.
@@
 Heuristics $n$ $d$ Ran-Greedy Deg-Greedy OnlineMIS ReduMIS 30000 10 7170.50 8951.75 9486.00 9505.38 30 3429.38 4464.12 4832.88 4914.88 100 1386.00 1817.25 1963.12 2012.62 300 570.50 738.00 794.75 815.50 1000 205.63 260.75 279.88 287.50 3000 106.00 106.12 100000 10 23990.00 29856.38 31613.38 31650.62 30 11444.38 14870.12 16128.88 16287.50 100 4615.25 6064.50 6564.00 6702.38 300 1886.62 2470.75 2661.50 2714.75 1000 83.50 941.38 959.75 300000 10 71954.38 89487.12 94622.88 94799.75 30 34318.88 44653.38 48234.38 48713.00 100 13850.12 18190.38 19676.50 20061.38 300 5688.00 7987.50 8141.25 1000 2831.38 3000 802.00 1000000 10 239749.12 312462.62 315630.88 30 114397.00 158625.29 161622.75 100 46161.12 64915.00 66539.75 300 26458.75 1000 6890.00 9473.00 3000000 10 719348.75 904613.12 942475.57 30 343479.38 459671.29 479938.75
-C.2 Serialization
-Figures 4 and 5 shows the percentage to choose the smallest possible degree node on different part of the serialization across various algorithms for all ER graphs and all BA graphs with nodes $n\leq 3000$, respectively.
+C.3 Comparison between Deg-Greedy and LTFT
+We conducted the exact same analysis as in Section 5.1 for BA graphs. The results are in Figure 2.
+Figure 2: Percentage of rounds when LTFT selects the node with smallest possible degree on BA graphs.
+On BA graphs, the percentage to choose the smallest possible node is generally lower, but on denser BA graphs LTFT also behaves more like degree-based greedy.
+C.4 Serialization
+Figures 3 and 4 shows the percentage to choose the smallest possible degree node on different part of the serialization across various algorithms for all ER graphs and all BA graphs with nodes $n\leq 3000$, respectively.
 The serialization process is discussed in Section 5.2.
-Missing bars are algorithms which we do not get results due to computational limit, same as in Tables 1 and 2.
+Missing bars are algorithms which we do not get results due to computational limit, same as in Tables 1 and 5.
 Figure 3: The percentage to choose the smallest possible degree node on different part of the serialization for all ER graphs It reinforces our observations in Section 5.2.
@@
 They have an easily found large MIS, which is the $m$ nodes initially in the graph at the start of the BA generation process.
-(Albert & Barabási, 2002).
+[46].
 From Table 5, we can see many algorithms can find these MIS and report a MIS size of $m$ for these graphs.
 This suggests that for this special type of BA graphs, our serialization analysis can observe different characteristics from other BA graphs.
-C.3 Comparison between Deg-Greedy and LwD
+C.5 Comparison between Deg-Greedy and LwD
 LwD, similar to LTFT, is also a non-backtracking MDP based algorithm which picks nodes sequentially.
@@
 From the heatmaps for different 1/3 parts, we can see the percentage increases from the 1st 1/3 to the 3rd 1/3 for all the datasets (different parameters of the synthetic graphs).
-This aligns with our “counterfactual” serialization results for LwD in Section 4.2, where we also observe the percentage increases clearly from 1st 1/3 to 3rd 1/3.
+This aligns with our “counterfactual" serialization results for LwD in Section 5.2, where we also observe the percentage increases clearly from 1st 1/3 to 3rd 1/3.
 This shows that our serialization method in Section 5.2 can reflect the pattern correctly.
-The percentages here in the heatmap is smaller than the percentages for “counterfactual” serialization in the bar graphs in Section 4.2.
-This is likely due to the fact that in the “counterfactual” serialization we repeat the serialization process for $100$ times and report the highest percentages we get, while here we only do serialization once for each step.
-C.4 Local search
+The percentages here in the heatmap is smaller than the percentages for “counterfactual" serialization in the bar graphs in Section 5.2.
+This is likely due to the fact that in the “counterfactual" serialization we repeat the serialization process for $100$ times and report the highest percentages we get, while here we only do serialization once for each step.
+C.6 Local search
 Table 9: Adding local search as a post-processing procedure. This is the full graph for local search described in Section 5.3.
-Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER Graphs 100,10 29.50 (0.25) 30.50 30.50 30.62 (0) 27.00 (2) 30.38 (0) 29.00 (0.38) 30.25 (0) 100,30 13.62 (0) 14.00 14.75 14.50 (0) 12.00 (0.50) 14.38 (0) 13.25 (0.13) 13.88 (0) 300,10 92.12 (0.50) 93.88 94.38 94.75 (0) 91.50 (0.88) 94.25 (0) 90.12 (1.50) 93.50 (0) 300,30 44.75 (0.25) 47.88 47.88 47.62 (0) 43.00 (1.25) 46.88 (0) 44.00 (0.75) 44.62 (0.74) 300,100 16.12 (0) 18.00 18.38 18.00 (0) 14.25 (0.87) 17.12 (0.12) 16.25 (0) 16.88 (0.26) 1000,10 305.38 (2.13) 314.75 316.13 315.62 (0) 308.12 (3) 311.25 (0) 300.00 (3.00) 306.38 (2.50) 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 148.50 (5.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 59.00 (3.38) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 1000,300 22.25 (0) 25.00 25.75 24.62 (0) 21.25 (1.25) 20.75 (1.63) 22.62 (0) 21.62 (0.74) 3000,10 913.62 (6.50) 947.25 954.25 950.88 (0) 927.00 (8.00) 935.38 (1.26) 900.62 (12.37) 918.00 (16) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 453.88 (11.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) 3000,100 185.62 (2) 194.38 201.50 200.38 (0) 180.25 (9.87) 191.50 (0.75) 185.62 (1.62) 182.25 (10.87) 3000,300 74.00 (0.50) 77.63 80.75 78.88 (0) 68.00 (4.50) 74.38 (0.50) 3000,1000 23.38 (0) 26.00 26.25 21.38 (1.50) 23.62 (0) BA Graphs 100,5 39.25 (0) 39.50 39.50 39.50 (0) 36.62 (2.74) 39.50 (0) 38.50 (0.38) 39.38 (0) 100,15 21.00 (0) 21.63 21.63 21.62 (0) 18.88 (0.50) 21.62 (0) 20.62 (0) 21.25 (0) 300,5 122.38 (0.26) 123.13 123.13 123.12 (0) 120.25 (2.63) 123.12 (0) 118.62 (3.00) 123.00 (0.12) 300,15 70.00 (0.75) 71.38 71.38 71.38 (0) 67.62 (2.00) 70.75 (0) 66.62 (1.87) 70.00 (0.50) 300,50 39.25 (1.75) 49.88 50.00 50.00 (0) 32.38 (5.13) 50.00 (0) 43.62 (1.87) 50.00 (0) 1000,5 413.38 (2) 417.13 417.13 417.12 (0) 412.62 (1.74) 416.00 (0) 400.25 (14.50) 417.12 (0) 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 235.50 (4.50) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 110.25 (2.37) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 1000,150 90.88 (7.63) 150.00 150.00 150.00 (0) 112.75 (3.13) 150.00 (0) 87.62 (5.37) 3000,5 1241.50 (4.75) 1257.00 1257.13 1255.62 (0) 1242.12 (3.37) 1248.50 (0.38) 1213.12 (35.87) 1254.75 (0.37) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 723.50 (12.88) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37) 3000,50 339.00 (3.62) 362.63 369.75 368.25 (0) 348.50 (9.62) 341.00 (4.12) 334.00 (10.12) 3000,150 142.88 (2.38) 160.25 165.75 164.00 (0) 150.50 (6.50) 146.50 (1.75) 3000,500 172.12 (9.12) 500.00 500.00 500.00 (0) 229.62 (5.74)
-Table 8 shows the full results after incorporating $2$-improvement local search from the ARW local search algorithm (Andrade et al., 2012) as a post-processing step, which is discussed in Section 4.3.
-C.5 More results on the ratio
-In addition to what we show in Section 4.4, Figure 7 shows the ratio of MIS size to $\frac{n\ln d}{d}$ for ER graphs with number of nodes $n$ and average degree $d$ across more algorithms.
-We can see that ReduMIS, OnlineMIS, and iSCO has consitently high ratios more than $1.2$.
-Ran-Greedy stays around $1.0$ for all $(n,d)$.
-Other algorithms, including Deg-Greedy, all have higher ratios for sparser graphs, but lower ratios (close to $1$) for denser graphs.
+Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER Graphs 100,10 29.50 (0.25) 30.50 30.50 30.62 (0) 30.63 (0) 30.38 (0) 29.00 (0.38) 30.25 (0) 100,30 13.62 (0) 14.00 14.75 14.50 (0) 14.50 (0) 14.38 (0) 13.25 (0.13) 13.88 (0) 300,10 92.12 (0.50) 93.88 94.38 94.75 (0) 94.63 (0) 94.25 (0) 90.12 (1.50) 93.50 (0) 300,30 44.75 (0.25) 47.88 47.88 47.62 (0) 47.63 (0) 46.88 (0) 44.00 (0.75) 44.62 (0.74) 300,100 16.12 (0) 18.00 18.38 18.00 (0) 18.00 (0) 17.12 (0.12) 16.25 (0) 16.88 (0.26) 1000,10 305.38 (2.13) 314.75 316.13 315.62 (0) 310.75 (0.75) 311.25 (0) 300.00 (3.00) 306.38 (2.50) 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 159.13 (0.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 60.88 (0.75) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 1000,300 22.25 (0) 25.00 25.75 24.62 (0) 23.25 (0) 20.75 (1.63) 22.62 (0) 21.62 (0.74) 3000,10 913.62 (6.50) 947.25 954.25 950.88 (0) 927.38 (4.13) 935.38 (1.26) 900.62 (12.37) 918.00 (16) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 469.75 (5.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) 3000,100 185.62 (2) 194.38 201.50 200.38 (0) 189.75 (4.12) 191.50 (0.75) 185.62 (1.62) 182.25 (10.87) 3000,300 74.00 (0.50) 77.63 80.75 78.88 (0) 70.88 (1.63) 74.38 (0.50) 3000,1000 23.38 (0) 26.00 26.25 23.00 (0) 23.62 (0) BA Graphs 100,5 39.25 (0) 39.50 39.50 39.50 (0) 39.50 (0) 39.50 (0) 38.50 (0.38) 39.38 (0) 100,15 21.00 (0) 21.63 21.63 21.62 (0) 21.63 (0) 21.62 (0) 20.62 (0) 21.25 (0) 300,5 122.38 (0.26) 123.13 123.13 123.12 (0) 123.00 (0) 123.12 (0) 118.62 (3.00) 123.00 (0.12) 300,15 70.00 (0.75) 71.38 71.38 71.38 (0) 71.25 (0) 70.75 (0) 66.62 (1.87) 70.00 (0.50) 300,50 39.25 (1.75) 49.88 50.00 50.00 (0) 50.00 (0) 50.00 (0) 43.62 (1.87) 50.00 (0) 1000,5 413.38 (2) 417.13 417.13 417.12 (0) 415.88 (0.13) 416.00 (0) 400.25 (14.50) 417.12 (0) 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 242.12 (0.12) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 114.00 (0.50) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 1000,150 90.88 (7.63) 150.00 150.00 150.00 (0) 150.00 (0) 150.00 (0) 87.62 (5.37) 3000,5 1241.50 (4.75) 1257.00 1257.13 1255.62 (0) 1245.38 (2.38) 1248.50 (0.38) 1213.12 (35.87) 1254.75 (0.37) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 729.38 (7.13) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37) 3000,50 339.00 (3.62) 362.63 369.75 368.25 (0) 359.25 (2.13) 341.00 (4.12) 334.00 (10.12) 3000,150 142.88 (2.38) 160.25 165.75 164.00 (0) 155.38 (2.63) 146.50 (1.75) 3000,500 172.12 (9.12) 500.00 500.00 491.00 (0) 229.62 (5.74)
+Table 9 shows the full results after incorporating $2$-improvement local search from the ARW local search algorithm [33] as a post-processing step, which is discussed in Section 5.3.
+Appendix D Perspective for theoreticians: Does KaMIS refutes famous conjecture?
+For theoretical analysis for MIS on ER graphs and regular graphs, see [51, 84, 54, 49].
+In ER graphs with $n$ nodes and average degree $d$, the MIS has size $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$, and simplest random greedy achieves half-optimal at $\frac{n\ln d}{d}$ [52].
+Yet, there is no known polynomial-time algorithm which can achieve MIS size $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ and it is conjectured that polynomial-time algorithms cannot do better than $(1+o(1))\frac{n\ln d}{d}$ [51].
+Thus it is natural to measure the goodness of an algorithm by the ratio of the MIS size obtained to $\frac{n\ln d}{d}$.
+This ratio can faciliate comparison across different $n$ and $d$’s.
+Figure 6 shows the ratio of MIS size to $\frac{n\ln d}{d}$ for ER graphs with number of nodes $n$ and average degree $d$ across various algorithms.
+Surprisingly for ReduMIS and OnlineMIS this ratio is consistently larger than $1.2$ and often larger than $1.3$ even for fairly large $n,d$.
+These results seem to bring the conjecture in doubt, but not disprove per se since the conjecture concerns asymptotically large $n$ and $d$.
+But our results could encourage further analysis and potential collaboration between researchers on theoretical and empirical aspects of MIS on random graphs.
 Figure 6: Heatmap for ratios of MIS size to $nln(d)/d$ on ER graphs with number of nodes $n$ and average degree $d$.